### v2

# Negative Sampling in Recommendation: A Survey and Future Directions

Abstract.
Recommender system (RS) aims to capture personalized preferences from massive user behaviors, making them pivotal in the era of information explosion. However, the presence of “information cocoons”, interaction sparsity, cold-start problem and feedback loops inherent in RS make users interact with a limited number of items. Conventional recommendation algorithms typically focus on the positive historical behaviors, while neglecting the essential role of negative feedback in user preference understanding. As a promising but easy-to-ignored area, negative sampling is proficients in revealing the genuine negative aspect inherent in user behaviors, emerging as an inescapable procedure in RS. In this survey, we first discuss existing user feedback, the critical role of negative sampling and the optimization objectives in RS and thoroughly analyze challenges that consistently impede its progress. Then, we conduct an extensive literature review on the existing negative sampling strategies in RS and classify them into five categories with their discrepant techniques. Finally, we detail the insights of the tailored negative sampling strategies in diverse RS scenarios and outline an overview of the prospective research directions toward which the community may engage and benefit.
1. Introduction
Recommender systems (RS) have emerged as an effective solution to the information overloading issue, capable of capturing user preference from the massive behaviors and provide appropriate items to each user. Presently, the extensive deployment of recommendation algorithms underscores their transformative power, shaping the way people engage and navigate the choices with digital content. These encompass electronic commerce platforms (Taobao111https://www.taobao.com/, and Amazon222https://www.amazon.com/), social network (WeChat333https://weixin.qq.com/, and Facebook444https://www.facebook.com), lifestyle applications (Meituan555https://www.meituan.com/ and Google Maps666https://maps.google.com) and so on. These significant achievements across multiple platforms and applications underscore the continued significance of RS.
Practical RS frequently consists of more than millions of users and items, each characterized by hundreds of auxiliary attributes. Despite this, the existence of personalized preference, “Information Cocoons” and the unique feedback loops intrinsic to the RS lead to a scenario in which users only interact with a limited number of warm items, rendering the majority of items of the corpus remain consistently unfamiliar to them (Li et al., 2022; Lou et al., 2019; Chen et al., 2021c; Meng et al., 2020) (as shown in the left part of Fig.1). The core of RS lies in its ability to infer unobserved preferences from the available interaction data, with a prevailing understanding that their efficacy excels as the richness of interaction data amplifies. However, the interaction data is observational rather than experimental. Thus, the ubiquity of noise within the incomplete users’ historical interactions prevalent in real-world RS may lead to the sub-optimal decision-making in recommendation (Chen et al., 2023a; Gao et al., 2023a).
In alignment with other supervised tasks, providing appropriate supervised signals (both positive and negative feedback) is indispensable in the training phase of recommendation algorithms. The absence of either of these feedback types can inevitably give rise to bias in model training (Ma et al., 2023; Zhu et al., 2022; Chen et al., 2023b). However, explicit feedback remains conspicuously absent within most recommendation datasets. Even when encompassed within datasets of such information, certain studies (Zhang et al., 2013; Ding et al., 2020; Chen et al., 2023b) have unearthed that these negative feedback indeed mirror user preferences to some extent. For instance, the behavior of assigning a low rating to a particular movie may paradoxically signify that the user has already viewed the movie, as this selection from an extensive movie pool subtly reveals the user’s underlying predilections (Ding et al., 2023; Huang et al., 2023). Consequently, the integral exploration of negative sampling strategies that are proficient in unveiling genuine negative aspects inherent in user preferences emerges as an inescapable procedure in RS.
As illustrated in Fig. 1, the training stage of RS typically collect the real behaviors and select the potential negative samples (NS) via diverse sampling strategies simultaneously, and then optimize the recommender algorithm on these given positive and negative signals. The profound impact of distinct negative sampling strategies on the performance of recommendation algorithms in disparate recommendation scenarios underscores the intricacies inherent in this critical domain. Moreover, existing academic endeavors pertaining to negative sampling algorithms are diverse in RS. Varied resolutions are observed in addressing NS across distinct recommendation scenarios, indicating the absence of a universal algorithm capable of concurrently tackling all recommendation tasks (Cheng et al., 2021; Li et al., 2021; Liu et al., 2021b; Zhao et al., 2019; Shi et al., 2023; Gong et al., 2022; Zhang et al., 2013; Kim et al., 2021a; Yang et al., 2022a). Finally, the predominant strategies for tackling matching and ranking tasks revolve around the random sampling strategies (Zhao et al., 2016) and the meticulous integration of genuine negative feedback (Lian et al., 2020a; Shen et al., 2021; Gu et al., 2021) for the industrial recommender systems. Currently, there lacks a comprehensive survey that systematically categorizes and presents the existing studies. Such a survey can serve as a handbook to offer a unified exploration of the varied landscape surrounding negative sampling algorithms in RS, thereby addressing the aforementioned issues.
Given the significance and popularity of RS, there are several recently published surveys also reviewed this area from different perspectives, including debiased recommendation (Chen et al., 2023a), cross-domain recommendation (Zang et al., 2022), graph-based recommendation (Gao et al., 2023c), knowledge graph-based recommendation (Guo et al., 2020b), causal inference-based recommendation (Gao et al., 2024), large language model-based recommendation (Fan, 2024; Wu et al., 2024c) and evaluation in recommendation (Bauer et al., 2023). As to negative sampling, there are surveys in other domains, including knowledge graph completion (Cai et al., 2023), deep metric learning (Kaya and Bilge, 2019), neural semantic retrieval (Guo et al., 2022; Chen et al., 2022c), image retrieval (Chen et al., 2022c), and representation learning on various types of data (Liu and Tang, 2021; Qian et al., 2020; Xu et al., 2022; Yang et al., 2020b; Chen et al., 2020a; Kamigaito and Hayashi, 2022; Wang et al., 2021a; Khoshraftar and An, 2022; Wang et al., 2017a). Compared to these research domains, negative sampling techniques employed in RS possess distinctive characteristics, necessitating a comprehensive review and summary. To bridge this gap, we review the negative sampling strategies in RS, classify them into five categories with the discrepant techniques and detail the insights of the tailored negative sampling strategies in diverse recommendation scenarios.
This survey encompasses over 230 papers from the top conferences and journals within the field of recommender systems, which focus on the analysis and the design of novel negative sampling techniques. We provide a comprehensive collection of negative sampling strategies in RS from various perspectives. In particular, we search related works from the top-tier conferences on the past 20 years, including SIGIR, WWW, KDD, ACL, AAAI, IJCAI, ICDM, ICDE, RecSys, WSDM, and CIKM, as well as the leading journals such as TOIS, TKDE, TORS, TKDD, KBS, etc, with the keywords “recommendation”, “collaborative filtering”, “ranking”, “re-ranking” and “matching” in conjunction with “negative samples”, “re-weighting” and “sampling”. At the same time, to prevent omissions of relevant literature, we utilize Google Scholar, DBLP, and Arxiv to search for related studies. After getting the list of papers, we also check their references to retain the unobserved papers about negative sampling strategies. We provide Figure 2 to illustrate the statistics of our collected publications related to negative sampling in recommendation with their publication year and venue. To facilitate this survey’s role as a handbook for negative sampling techniques in RS, we will release a GitHub repository to encompass all reviewed papers along with their corresponding codes at the following link: https://github.com/hulkima/NS4RS.
We believe this survey is beneficial for those following academic researchers and industry developers in RS: 1) who are a new to the negative sampling problem and seek a handbook with an overview of related works; 2) who are aspiring to delve into the investigation of negative sampling algorithms but are unsure of where to begin; 3) who are grappling with efficiency or performance issues in their recommendation algorithms and seeking solutions through negative sampling strategies. We summarize the contribution of this survey as follows.
We unify the achievements made in the communities of negative sampling and RS by presenting a comprehensive review of existing related studies, which is the pioneering survey of this promising but easy-to-ignored area.
We summarize the fundamental challenges of negative sampling in practical RS from three aspects, delineate existing research works into five categories regarding representative methods and characteristics.
We detail the effectiveness mechanism and meaningful insights of the tailored negative sampling methods in diverse recommendation scenarios and tasks.
We outline further potential research directions for negative sampling in recommendation by intensively analyzing the most advanced technologies in recommendation and the fundamental challenges of negative sampling.
2. Necessity and Challenges of Negative sampling
Negative sampling emerge as a critical and irreplaceable element in recommendation. Within this specific domain, how to effectively discover users’ real negative interests in negative sampling stands as a central research objective (Chen et al., 2023c). In this section, our first introduce two types of feedback within recommendation, then we highlight the significance of negative sampling algorithms in recommendation, with particular attention to the crucial role played by hard negative samples, and finally we analyze the existing challenges of negative sampling techniques in recommendation.
2.1. Introduction of User Feedback in Recommendation
Constrained by the exposure bias and popularity bias inherent in the real-world RS, user interaction reveals a notable unbalancedness, which constitutes a computationally insurmountable problem in recommendation (Huang et al., 2023). Regarding the available user feedback, both positive and negative feedback can reflect the user’s opinions on diverse items in the corpus, which are equally essential for modeling user preferences in RS (Zhao et al., 2018; Wu et al., 2022b). Leveraging such data to its maximum capacity can provide the recommender with a significant information gain. Following the classical studies (Wang et al., 2021c; Xu et al., 2021; Wu et al., 2022a), user feedback can be categorized into two different types as explicit feedback and implicit feedback 777Apart from explicit feedback and implicit feedback, the presence of real negative feedback is a notable factor in practical RS (Xie et al., 2021). Due to the scarcity of this type of feedback, we discuss their definitions and applications in Sec .4.4.:
Explicit Feedback: Explicit Feedback materializes in the form of precise ratings to provides unequivocal insights into user preferences (e.g., from 1.0 to 5.0).
Implicit Feedback: Implicit Feedback encompasses data derived from user behaviors without any ratings and reviews (e.g., residence time and whether to choose dislike).
From the above two types of feedback, implicit feedback is fundamentally different from the explicit one, which is captured in a structured manner (Saito et al., 2020a). Precisely, explicit feedback delivers direct insights into user preferences, while implicit feedback unravels concealed user interests. Regarding the explicit feedback, existing RS algorithms (Sun et al., 2021; Li et al., 2023; Dai et al., 2023) typically regard the user feedback with lower value (e.g., ratings of [1.0, 2.0, 3.0]) as the negative signals, which may not necessarily reflect the users’ real preferences due to the existence of selection bias in recommendation (Chen et al., 2023a). Even for the item explicitly rated as 1.0, they have been chosen by this user from an extensive candidate pool. Therefore, such negative feedback may not be as too ”negative” as they reflect this user’s preference to some extent. That is, these items appear more ”negative” than those the user is interested in, yet somewhat more “positive” when compared to items the user has never interacted with. As noted in (Wang et al., 2021c), implicit feedback can be classified into true positives, false positives, true negatives, and false negatives, among which this survey primarily concentrates on the negative aspect. For the negative implicit feedback, traditional RS methods configure all of the uninteracted items as the negative sample pool, which lacks explicitness. Meanwhile, it necessitates the recommender to elucidate users’ preferences, which may introduce long-tailed information and exposure bias in RS.
2.2. Role of Negative Sampling in Recommendation
Real-world recommender systems frequently involve more than millions of users and items, rendering the integration of all corpus into the training process prohibitively expensive (Zhang et al., 2013; Shi et al., 2023; Zhou et al., 2020). Despite the presence of a substantial number of users and items, the existence of “Information Cocoons” and the unique feedback loops within recommendation result in the majority of users interacting with only a limited number of items in the corpus (Li et al., 2022; Lou et al., 2019; Chen et al., 2021c). As a consequence, this gives rise to the well-documented issue of data sparsity in recommendations (Ma et al., 2024a, 2023; Cao et al., 2022; Xie et al., 2022b). The dynamic nature of user interests also necessitates a continuous process of adaptation and updating for recommenders (Zhao et al., 2016; Xu et al., 2021; He et al., 2016). Furthermore, the perpetual introduction of new users and items (both devoid of an adequate reservoir of historical interaction data) engenders the notorious cold start problem (Wei et al., 2021; Togashi et al., 2021).
Large-scale recommendation applications typically encompass multiple phases to narrow down the candidate items for the recommender, thereby alleviating the aforementioned issues and significantly reducing their computational complexity (Huang et al., 2023; Ding et al., 2023). These algorithms generally require both positive and negative examples to model the users’ personalized preferences. Negative sampling is the critical and irreplaceable element in recommendation that could potentially improve the modeling of dynamic user preferences with their sparse interactions. Its crucial secret lies in its ability to select samples from each user’s vast collection of unobserved items, specifically tailored to enhance the model’s optimization within its current state (Shi et al., 2023; Chen et al., 2023b). To achieve this, it necessitates the deliberate selection of NS characterized by (1) heightened informativeness, (2) increased discriminative capacity, and (3) enhanced precision. Such samples are typically termed as hard negative samples (HNS), which are derived from their tendency to encompass more comprehensive information compared to the randomly selected samples from the corpus. Incorporating them into the training process serves to balance the positive and negative information within the dataset, thereby ensuring an unbiased optimization of the recommender. Drawing from this foundational definition of HNS, their gradient directions and magnitudes tend to vary from other random NS. Theoretically, more HNS can not only expedite the recommender’s convergence but also rectify the optimization direction of the global gradient, thus making it computationally possible to predict the appropriate items for users from the sparse implicit interactions (Ma et al., 2023; Zhang et al., 2021a). In the practical deployment, the inclusion of HNS enables recommenders to elevate their effectiveness and robustness simultaneously, which has been verified in many relevant studies.
2.3. Optimization Objective in Recommendation
How to design a effective and robust loss function is the universal but essential research topic in recommendation, which is bidirectionally coupled with the negative sampling techniques in recommendation to jointly establish recommenders’ optimization objective and generalization ability. Here, the loss function defines the optimization objectives for positive and negative samples, with different losses exhibiting varying degrees of dependency on and utilization of negative samples. In turn, the quality, quantity, and distribution of negative samples directly influence the gradient direction and the convergence of loss function, potentially undermining its effectiveness. Therefore, a series of studies in recommendation start to investigate effective loss functions to optimize recommenders’ convergence towards the global minima. Regarding the ratio between positive and negative samples, these approaches can be broadly categorized into One2One (one positive to one negative) and One2More (one positive to multiple negatives) schemes.
One of the most representative One2One optimization objectives in recommendation is BPR loss (Rendle et al., 2009), which assumes that user prefer the positive item over the non-observed samples. Given user , positive item and the sampled negative instance , BPR loss can be formulated as , where denotes whole training triples, denotes the Sigmoid function, and represent the score predicted by the recommender. It maximizes the score difference between positive and negative samples, which forces the recommender with relative item preferences rather than absolute scores , thereby closely aligning the goal of top-k recommendation tasks. Compared to the One2One optimization objective, One2More has attracted increasing attention due to its ability to accelerate recommender optimization and mitigate the impact of potential false negative samples on the global gradient by incorporating several negative samples (Wu et al., 2024b; Ma et al., 2023). For instance, BSL (Wu et al., 2024a) applies the same Log-Expectation-Exp structure to positive examples and multiple negatives to maintain the robustness of recommenders to the noisy positives. Adap- (Chen et al., 2023e) defines the Adaption and Fine-grained principles and utilizes Superloss to adaptively adjust the temperature by monitoring the sample losses for each user. S-DPO (Chen et al., 2024) instills ranking knowledge into the LLM via pairing multiple negatives with the positive sample to help LLM-enhanced recommenders distinguish preferred items from negatives. Benefiting from the CL paradigm, AdvInfoNCE (Zhang et al., 2023b) and PCL (Tang et al., 2024b) adaptively assigns hardness to each negative instance in an adversarial manner and collaboratively combine the absolute constraints of BCE/CE loss with CL to make the representations uniformly distributed respectively, thereby mining informative knowledge from the hard negatives. The discussion of optimization objectives in recommendation further underscores the critical role of negative sampling. We believe that the negative sampling strategies summarized in this survey can be seamlessly integrated with existing optimization objectives and will benefit from future advances in this technique to drive progress in recommendation.
2.4. Challenges of Negative Sampling in Recommendation
Real-world recommendation applications still face fundamental challenges with the false negative problem, the trade-off among accuracy, efficiency, and stability, and the universality across different tasks, goals, and datasets.
How to precisely identify negative feedback with the substantial knowledge in recommender systems?
How to balance the accuracy, efficiency and stability in model training?
How to handle multiple scenarios with diverse objectives and data availability in recommendation?
2.4.1. False Negative Problem
Traditional recommenders randomly select items that users have not interacted with as negative samples (Rendle et al., 2009; Liu et al., 2021a; Zhao et al., 2018). In addition to these potentially meaningless random samples, some recommendation algorithms attempt to sample informative yet not excessively challenging negative items from the random candidates (Zhang et al., 2013; Shi et al., 2023; Chen et al., 2023b). These samples compel the recommender to delve deeper into modeling the boundary between positive and negative samples in many complex scenarios, thereby improving the recommender’s ability to discriminate among the unobserved items. Some recent negative sampling methods (Shi et al., 2023; Ma et al., 2023; Liu et al., 2023a) have explicitly distinguished these items as the HNS and further given the conception of false negative samples (FNS) in recommendation. We also leverage Fig. 3 to illustrate the distribution and the toy example of each type of samples. For instance, radios and compact discs that share the same category of “electronics” may be false negative samples (potential positive samples). Conversely, watches, tall glasses, and high-heeled shoes can serve as the negative samples with the “hardness” decreasing, which exhibit less associated with the category of “electronics”. The corresponding definitions of these samples can be presented as:
Negative Samples (NS): As discussed in Sec. 2.1, user feedback can be categorized as explicit feedback and implicit feedback. Regarding the former, NS typically denote the items which are rated with low values (e.g., [1, 2, 3]). While for the latter, NS commonly refer the complete set of all items in the corpus that user has not interacted with. This type of samples typically serves as the fundamental candidate of all existing negative sampling strategies.
Hard Negative Samples (HNS): The samples that possess more information than the majority of NS, such as those associated with user ’s positive items or items engaged with user ’s social connections.
False Negative Samples (FNS): The samples that are erroneously identified as negative samples and subsequently fed into the recommender optimization, which correspond to the real interest of user .
As highlighted in the discussion of exposure bias in (Chen et al., 2023a), user behaviors are collected from the observed data, which makes them inherently noisy, stochastic, and incomplete. However, the subsequent negative sampling strategies have generally focused on augmenting the ”hardness” of HNS. This includes but is not limited to introducing supplementary information (Sun et al., 2015; Wang et al., 2021b; Zhao et al., 2019; Chen et al., 2023b), enhancing the sampling weight of HNS (Zhang et al., 2013; Shi et al., 2023; Ma et al., 2023), and employing positive representation interpolation (Huang et al., 2021; Shi et al., 2022). In this case, certain NS considered overly ”hard” by these negative samplers might in fact represent the potential positives (Chen et al., 2023a). Therefore, incorporating them as the negative samples into the recommender optimization may introduce bias during the back-propagation process, potentially resulting in the poor convergence speed and quickly falling into local optimum and ultimately leading to the False Negative Problem in recommendation. The ”hardness” of samples varies significantly across different recommendation algorithms and datasets, resulting in substantial instability in parameter selection (Zhu et al., 2022; Ding et al., 2020; Shi et al., 2023; Ma et al., 2023) and further aggravating this problem. While some prior works (Saito, 2020; Saito et al., 2020b; Li et al., 2024; Ha et al., 2024) attempt to alleviate this phenomenon by modeling the unbiased representation with the Missing-Not-At-Random feedback, their efforts have largely centered on the theoretical guidance or the refinement of learning objectives, neglecting the more straightforward and effective strategy of negative sampling. To summarize, the false negative problem stands as a prevalent and urgent challenge of this specific research domain to be solved in RS.
| Algorithms | Venue | Year | Scenarios | Sampling criteria |
| RNS (Rendle et al., 2009) | UAI | 2009 | CF | Random sampling | ||||||
| DNS (Zhang et al., 2013) | SIGIR | 2013 | CF | Top-ranked | ||||||
| PNS (Rendle and Freudenthaler, 2014) | WSDM | 2014 | CF | Popularity | ||||||
| SRNS (Ding et al., 2020) | NeuIPS | 2020 | CF | Variance-based | ||||||
| MixGCF (Huang et al., 2021) | KDD | 2021 | GR | Interpolation | ||||||
| DNS* (Shi et al., 2023) | WWW | 2023 | CF | Highest-score scope | ||||||
| RealHNS (Ma et al., 2023) | Recsys | 2023 | CDR |
2.4.2. Trade-off among Accuracy, Efficiency and Stability
The ”information overload” phenomenon has compelled recommenders to grapple with the issue of striking an appropriate trade-off among accuracy, efficiency, and stability (Chen et al., 2023a; Gao et al., 2024). Negative sampling is no exception to this challenge (Ding et al., 2020; Li et al., 2021; Wang et al., 2020). Accuracy pertains to the precision of model’s predictions regarding user preferences, which is critical in evaluating the performance of recommenders. Efficiency focuses on how many queries the recommender can process within a unit of time, directly influencing its scalability and user experience. And stability concerns the consistent performance of the recommender across diverse datasets.
Nonetheless, these three aspects are generally interrelated in RS, making it challenging to balance them simultaneously. We furnish the sampling criteria, as well as the time complexities of candidate generation and negative selection for seven classic negative sampling algorithms, arranged in chronological order in Table 1. This illustrates that with the advancement in the accuracy of negative sampling techniques, there is a synchronous increase in their time complexity, underscoring the delicate trade-off between accuracy, efficiency and stability (Rendle et al., 2009; Rendle and Freudenthaler, 2014; Zhang et al., 2013; Ding et al., 2020; Shi et al., 2023; Huang et al., 2021; Ma et al., 2023). This is due to the fact that enhancing accuracy may lead to increased computational complexity, resulting in decreased efficiency. Some efficiency-improving optimization techniques may exhibit greater sensitivity to data variations and neglect specific details, potentially impacting the model’s accuracy and stability. Simultaneously, the pursuit of stability may increase time and space complexity, conflicting with the balance between accuracy and efficiency. Therefore, achieving the optimal balance among these three factors within negative sampling is a complex and challenging task.
2.4.3. Universality with Different Tasks, Objectives, and Datasets
In essence, the prevailing recommendation algorithms encompass different datasets with diverse distribution, distinct training objectives (e.g., matching and ranking), various recommendation scenarios (e.g., collaborative filtering, multi-modal recommendation, multi-behavior recommendation, sequence recommendation, and cross-domain recommendation), and diverse evaluation methods (random sampled testing, popularity-based sampled testing, and full testing). Although general negative sampling strategies (RNS (Rendle et al., 2009), PNS (Rendle and Freudenthaler, 2014), DNS (Zhang et al., 2013)) are available, some tailored negative sampling techniques designed for specific recommendation tasks can outperform these general methods (Cheng et al., 2021; Li et al., 2021; Liu et al., 2021b; Zhao et al., 2019; Shi et al., 2023; Gong et al., 2022; Zhang et al., 2013; Kim et al., 2021a; Yang et al., 2022a). Upon a comprehensive examination of the current negative sampling algorithms, it is evident that none of these approaches can perfectly match all recommendation tasks. In practical implementation, employing different negative sampling strategies for various tasks presents a challenge to the efficiency and stability of the recommender. This challenge underscores the obstacles that large-scale recommender systems must overcome when transitioning to the real-world online deployment.
3. Literature review of Negative Sampling in Recommendation
Let and be the number of users and items , we can define a user-item interaction matrix , where denotes the user has interacted with the item , and denotes has never interacted with . Negative sampling aims to select the appropriate instances from the unobserved item candidates as the negative samples to support the recommender optimization. We conduct a comprehensive literature review of existing negative sampling strategies in recommendation, classify them into five categories according to their technical trajectories, provide the overall ontology of our taxonomy in Fig. 4 and additionally offer five tables (from Table 2 to Table 6) to illustrate their core contributions, strengths, and drawbacks: 1) Static negative sampling typically samples negative instances with the static probability; 2) Dynamic negative sampling attempts to dynamically select negative instances with the pre-established sampling criterion; 3) Adversarial negative generation leverages the adversarial learning paradigm to sample or generate plausible items as the negative samples; 4) Importance re-weighting aims to identify the importance of each sample and assign diverse weights for them in a data-driven manner; and 5) Knowledge-enhanced negative sampling focuses on sampling negative instances with the mining of the implicit association from the auxiliary knowledge. In the following, we detail the effectiveness mechanisms and meaningful insights of these negative sampling strategies.
3.1. Static Negative Sampling
Traditional approaches in the initial stages of deep learning for recommendation relied on the static negative sampling (SNS) method whereby several items are selected from users’ unobserved items. The primary objective of SNS is to offer a diverse range of informative negative samples and facilitate the acquisition of more comprehensive user preference patterns by the recommender. We will introduce several types of SNS strategies (Uniform SNS, Predefined SNS, Popularity-based SNS, and Non-sampling SNS) in the following subsections.
3.1.1. Uniform SNS
Uniform SNS randomly selects items from the unobserved interactions of the user as the negative sample (Zhou et al., 2018; Dai et al., 2024a). This strategy is extensively employed across diverse recommendation scenarios (e.g., Collaborative-guided Recommendation (Rendle et al., 2009; He et al., 2017, 2023a; Wu et al., 2023b), Sequential Recommendaiton (Yang et al., 2022b; Zou et al., 2024), Social Recommendation (Yang et al., 2023c, 2024; Wu et al., 2023a, c), Multi-modal Recommendation (Wang et al., 2023a, 2024a), and Cross-domain Recommendation (Cao et al., 2022; Wang et al., 2024b)) and distinct recommendation applications (e.g., Point-of-Interest Recommendation (Lian et al., 2014), News Recommendation (Wu et al., 2019c), Song Recommendation (Tran et al., 2019) and API Recommendation (Wei et al., 2022)). Primarily, its ease-to-deploy mechanism makes it well-suited for large-scale RS, as it obviates the necessity for additional computations. It also introduces a degree of diversity, mitigating the risk of over-fitting and consequently enhancing recommender’s generalization and robustness. Moreover, this strategy facilitates the exploration of items that users have not interacted with, thereby aiding the recommender in uncovering novel recommendation opportunities and unexplored areas of user preference.
Notably, TDM (Zhu et al., 2018) denotes the positive sample as the leaf nodes that have interacted with the user and their ancestor nodes within the hierarchical tree structure, while randomly selecting nodes except positive ones in each level to constitute the set of negative samples. RSS (Petrov and Macdonald, 2022) also tries to propose a recency-based strategy to generate multiple positive samples out of a single user sequence simultaneously, which is orthogonal to negative sampling and can be applied independently. However, given that the primary focus of this survey lies in the examination of negative sampling in recommendation, there has been limited elaboration on the positive-augmented sampling methods.
3.1.2. Predefined SNS
Predefined SNS incorporates the pre-defined negative samples from the dataset into the recommenders’s training process (Song et al., 2015; Yu et al., 2018; Sun et al., 2021; Zhang et al., 2024b). For example, DRN (Zheng et al., 2018) employs real user behaviors (e.g., skipped, clicked, and ordered actions) within the dataset to delineate positive and negative samples. Additionally, a series of studies designate the ratings of in the explicit user interactions as positive samples. Within this context, ELLR (Song et al., 2015) and IF-BPR (Yu et al., 2018) segregate ratings of as negative samples, while SCoRe (Qin et al., 2020), ML-MGC (Sun et al., 2021) and many LLM-based recommenders (Lin et al., 2024; Xi et al., 2024; Dai et al., 2023; Li et al., 2023) extend this categorization to include ratings of as negative samples.
3.1.3. Popularity-based SNS
Relying on the assumption that the popularity of items may demonstrate users’ global preferences in the specific domain, popularity-based SNS typically selects negative samples based on the popularity of items, that is, the more popular the item is, the more likely it is to be selected as the negative sample (Gantner et al., 2012; Cheng et al., 2021; Ma et al., 2018; Rendle and Freudenthaler, 2014; Togashi et al., 2021; Wang et al., 2019; Li et al., 2018; He et al., 2016; Yang et al., 2023d). This encourages the recommender to incorporate more popular items as negative samples into the training process. For quick updating the recommender parameters with the new incoming data, eALS (He et al., 2016) assigns the weight of missing data based on the popularity of items by an incremental update strategy. SoftRec (Cheng et al., 2021) allows some unobserved items to have non-zero supervisory signals by generating soft targets with popularity distribution, thereby incorporating more unobserved knowledge into the training process. Similar to RSS (Petrov and Macdonald, 2022), SAE-NAD (Ma et al., 2018) adopts the general weighting scheme to assign higher confidence to each positive sample based on its frequency. In contrast, it assigns the weights of all negative examples to the same value (e.g., 1.0) for the Point-of-Interest recommendation task.
Notably, KGPL (Togashi et al., 2021) and MKR (Wang et al., 2019) employ the Popularity-based SNS on the graph structure. For instance, KGPL (Togashi et al., 2021) directly formulates the sampling confidence of negative samples based on their interaction frequencies across the entire dataset. Furthermore, it assigns reliable pseudo-labels to the interacted items of cold-start users by considering the reachable items under given meta-paths (i.e., the paths rooted at a user’s interacted items in KG). Different from the above studies, MKR (Wang et al., 2019) follows the negative sampling strategy of the classical work (Mikolov et al., 2013) in NLP, which reduces the sampling probability for the high-frequency items and augments it for the low-frequency items. This is motivated by the observation that less probable words tend to be more pivotal, while higher probability words such as ‘the’, ‘a’, and ‘to’ bear less weight in NLP tasks. It actually deviates from the core assumptions in recommendation, thus MKR adopts such a sampling strategy primarily to enhance the computational efficiency.
3.1.4. Non-sampling SNS
Non-sampling SNS considers the unobserved instances from the whole training data for recommender learning, thus avoiding negative sampling (Chen et al., 2020d, c, 2019d, b; Li et al., 2021). This strategy argues that negative sampling are highly sensitive to the data distribution and the number of negative samples, making them difficult to achieve the optimal performance in large-scale RS. To converge to a better optimum, EATNN (Chen et al., 2019d) first devises an efficient optimization method to learn from the whole training set without negative sampling, and reformulates a square loss function with rigorous mathematical reasoning for accelerating the training process in social-aware recommendation. ENMF (Chen et al., 2020c) stands as an extended edition of EATNN (Chen et al., 2019d) that extends its efficient learning method with the item-based and alternating-based forms, enabling its application for implicit feedback. Due to the behavior characteristic in multi-bahevior recommendation, EHCF (Chen et al., 2020d) links the prediction of each type of behavior in a transfer manner and applies non-sampling optimization for the recommender by considering all samples in each parameter update. Inspired by the memorization strategies, ENSFM (Chen et al., 2020b) employs the non-sampling technique to maintain the convergence into a better optimum in a stable way. JNSKR (Li et al., 2021) refine the existing non-sampling strategy into the modeling of the knowledge graph embeddings, which consists of entity-relation-entity triplets. Furthermore, it devises an efficient optimization approach to learn more precise entity embeddings and item representations within the knowledge graph, resulting in a stable enhancement of recommendation performance.
| Algorithms |
Advantages | Challenges | ||||||
| Uniform SNS (Rendle et al., 2009; Petrov and Macdonald, 2022; Wang et al., 2023a, 2024a; Zou et al., 2024) |
Accuracy | |||||||
| Predefined SNS (Song et al., 2015; Yu et al., 2018; Zheng et al., 2018; Lin et al., 2024; Xi et al., 2024; Dai et al., 2023; Li et al., 2023) | User-centric rating |
Data availability | ||||||
| Popularity-based SNS (Gantner et al., 2012; Cheng et al., 2021; Togashi et al., 2021; Wang et al., 2019; Li et al., 2018; He et al., 2016) |
| Non-sampling SNS (Chen et al., 2020d, c, 2019d, b; Li et al., 2021) |
Efficiency |
Summary: The in-depth investigation of behavior dependencies, advantages, and challenges of these above four SNS categories are documented in Table 2. Among them, Uniform SNS emerges as the most widely employed SNS method with its easy-to-deploy and time-efficient character. However, the randomness inherent in sampling introduces variability to its recommendation performance. Predefined SNS leverages authentic user-centric ratings for negative sample selection. Despite its straightforward implementation, it is critically reliant on the accessibility of user behaviors. In the incorporation of the mainstream preferences of the majority of users into the training process, Popularity-based SNS also introduces the inherent popularity bias and conformity bias. Non-sampling SNS incorporates the unobserved items of each user from the whole training data. Maintaining visibility over the entire corpus of the dataset enhances the recommender’s performance, yet encountering challenges in computational efficiency.
3.2. Dynamic Negative Sampling
To capture the informative negative samples in a universal manner, a series of Dynamic Negative Sampling (DNS) strategies has sprung up in recent years. DNS refers to selecting the item that is more relevant to the positive sample representations (or user representations) as the negative samples from the dynamically selected item candidates. For example, the pioneering dynamic negative sampling strategy (Zhang et al., 2013) generally selects the hardest item from the randomly selected item candidates. The subsequent work primarily focuses on modeling the relationship between users and items, or precisely tuning the sampling difficulty of the candidate set. Particularly, DNS have outperformed most of the existing negative sampling methods in different recommendation objectives and scenarios. We categorize existing DNS strategies into six groups: Universal DNS, User-similarity DNS, Knowledge-aware DNS, Distribution-based DNS, Interpolation DNS and Mixed DNS. In the subsequent parts, we will introduce their implementation details and effectiveness mechanisms in detecting informative negative samples and facilitating recommender optimization.
3.2.1. Universal DNS
Universal DNS selects the top-ranked item as the negative sample from the randomly selected item candidates (Wu et al., 2019b; Yang et al., 2020a; Shi et al., 2023; Ying et al., 2018; Brin and Page, 1998; Hao et al., 2021b; Ma et al., 2023, 2024a; Lai et al., 2024), which is proposed to select more informative negative samples with the current model state and has been widely applied in various recommendation scenarios.
Top-ranked universal negative selection. DNS (Zhang et al., 2013) hypothesizes that any unobserved item should not be ranked higher than any observed positive item and subsequently oversamples items top-ranked by the recommender from the randomly selected candidates. Due to the fact that the more popular an item is, the more times it acts as a positive sample in the conventional pairwise training procedure. DNS is expected to alleviate the popularity bias in RS. Along this line, a series of studies (Hu et al., 2018; Gong et al., 2022) enhance DNS from different perspectives. For example, MCRec (Hu et al., 2018) pretrains the nodes’ embedding with a pure matrix factorization method, computes the pairwise similarities between two consecutive nodes along a path instance, and then keeps top-K path instances with the highest average similarities. To efficiently identify hard negative instances, CA-FME (Li et al., 2020) samples negative instances from triple sides of the positive triplet, and grants larger sampling probability for harder negative samples according to their compatibility scores.
Dynamic negative selection with high-ranking scope. PinSage (Ying et al., 2018) argues that the highest-scored item is likely to be a potentially positive sample, leading to the false negative problem. Meanwhile, it also demonstrates that top-ranked items are more similar to positive items than randomly selected ones, which forces the recommender to learn to distinguish items at a finer granularity. Specifically, it utilizes the query items’ PageRank score to rank the uniformly selected candidates, randomly samples hard negative items from the items ranked between 2000 and 5000. ITSMGCN (Gong et al., 2022) proposes an negative sampler to detect the best sampling scheme by carefully tuning the number of item candidates and the ranking position of hard negatives in ranking diagram. DNS (Shi et al., 2023) further reduces the probability of selecting overly hard negative samples by expanding both the number of item candidates and the selection range in DNS.
Collaborative distillation of dynamic negative samples. To address the sparsity of positive feedback and the ambiguity of missing feedback in recommendation, CD-SG (Lee et al., 2019) follows the knowledge distillation (KD) paradigm and proposes collaborative distillation (CD). Specifically, the student model of CD-SG samples items with the dynamic soft target according to their ranking orders in the teacher model, with the aim to capture both positive and negative correlations and overcome the disadvantage of rank distillation that ignores negative feedback among items. Similarly, DE-RRD (Kang et al., 2020) reformulates the task of learning all the precise ranking orders in recommendation to a relaxed ranking matching problem within the framework of KD. It first defines both the detailed ranking orders among the interesting items and the relative ranking orders between the interesting items and the uninteresting items from the teacher’s recommendation list and then defines a relaxed permutation probability to locate all the interesting items higher than all the uninteresting items in student model, while maintaining the detailed ranking orders among the interesting items.
Dynamic negative selection with representation learning. With the advancement of representation learning in recommendation, a cohort of researchers have embarked on an investigation of the distinction between easy and hard negative samples from the perspective of representation learning. To illustrate, PDRec (Ma et al., 2024a) examines the inherent information disparities among observed instances (positive samples) and the latent relevance within the unobserved items (negative samples). It believes that following the emphasis on information-rich samples within the corpus as the positive instances, the deployment of hard negative sampling may concurrently exacerbate the data imbalance and over-fitting risks within recommenders. To address this issue, PDRec designs a noise-free negative sampling strategy to capture items with low confidences as the Safe Negative Samples (SNS) which exhibit no relevance to user preferences in model optimization. In contrast, RealHNS (Ma et al., 2023) meticulously investigates the definition of Real Hard Negative Samples (RHNS) (samples that are neither too easy nor too hard) and False Hard Negative Samples (FHNS) (samples that represents users’ actual preferences in the future) and then defines the equation: HNS=RHNS+FHNS. To effectively leverage these two concepts, RealHNS employs a curriculum learning framework to construct the multi-granular RHNS selector (coarse-grained candidates selector and fine-grained samples selector). This facilitates the dynamic sampling of high-quality RHNS that are suitable for recommender’s current state, thereby alleviating the issue of negative transfer.
3.2.2. User-similarity DNS
User-similarity DNS identifies similar users based on their historical behaviors and then dynamically selects negative samples according to this similarity association (Wu et al., 2019b; Wang et al., 2021b; Giobergia, 2022; Chen et al., 2019b). Some studies emphasize the explicit associations among users to approximate users’ conditional preference state, inferring the information dissemination path and collaborative relationships through the connections in social networks or explicit interaction behaviors. SamWalker (Chen et al., 2019b) proposes an social-based sampling strategy to both speed up the gradient estimation and reduce the time-consuming of the preference inferring from the massive unobserved feedback. To reduce the dependency on other users’ preferences, EvalRS (Giobergia, 2022) proposes a relatively audacious negative sampling strategy to capture neighboring users with their collaborative representations and regards item from the interaction list of the user’s neighbors that the user have not interacted as the negative sample. Conversely, other studies focus on implicit relationships, which refer to connections that are not directly observable. Such approaches aim to reveal the potential collaborative behaviors and interest similarities by mining users’ behavioral patterns. DEEMS (Wu et al., 2019b) believes that unobserved positive samples are more likely to exist correlations with observed positives than unobserved negatives. Based on this hypothesis, DEEMS designs a hedging gradient between the user- and item-centered correlations to rerank the unobserved positive samples before the negative ones. To handle the absence of explicit social information, SamWalker++ (Wang et al., 2021b) constructs the pseudo-social network, that is, similar users are connected with specific item nodes or community nodes. It proposes a subtle sampling strategy based on the random walk along the pseudo-social network to precisely model the triplet confidence and determine which data are used to update parameters.
3.2.3. Knowledge-aware DNS
Knowledge-aware DNS dynamically captures the informative samples with their attributes or other relevant features and selects the informative items related to the positive sample as negative samples (Lian et al., 2020b; Chen et al., 2023b, 2016; Wang et al., 2021d). This strategy is designed to incorporate task-specific information to intelligently select more relevant and informative negative examples for the specific task. In Location Recommendation, users’ negatively-preferred locations are typically mixed with potentially positive ones. However, informativeness differs from sample to sample, so treating them equally in optimization is far from optimal. GeoSAN (Lian et al., 2020b) designs geography-aware negative sampler to retrieve some nearest locations to the target location as candidates and employ negative sampling based on the uniform distribution or a popularity-based distribution. CITING (Chen et al., 2016) proposes a time-aware negative sampling strategy to select the tweets that were posted in close temporal proximity to his/her positive tweets as negative ones in Tweet Recommendation. RPRM (Wang et al., 2021d) also designs a novel negative sampling strategy to model the agreement of review properties between the users’ interacted items and the unseen items. To investigate the Debiased Recommendation, FairNeg (Chen et al., 2023b) dynamically perceives the group-level unfairness based on their performance disparity during the training process and adjusts each group’s corresponding negative sampling probability, intending to equalize all groups’ performance.
3.2.4. Distribution-based DNS
Distribution-based DNS analyzes the distribution pattern of positive and negative samples, then dynamically selects informative negative samples that do not overly disturb the training phase (Zhu et al., 2022; Ding et al., 2020; Lai et al., 2023). Some sampling strategies (Ding et al., 2020; Zhu et al., 2022) explore the statistical properties of both RHNS and FHNS to distinguish the relevant negative samples that are less informative or misleading. SRNS (Ding et al., 2020) discovers an empirical conclusion that only a limited number of samples are potentially crucial for the recommender, and false negative samples tend to get stable scores during training and further proposes a variance-based sampling function to distinguish negative samples. GDNS (Zhu et al., 2022) introduces an expectational gain-tuning sampler to leverage the high-gap expectation of users’ preference in the training process to measure the quality of negative samples and dynamically guide the negative sampling procedure. In contrast, DENS (Lai et al., 2023) argues that real-world interactions are primarily driven by certain relevant factors associated with the items while not encompassing all factors. Therefore, It delineates the relevant and irrelevant factors within items that are associated with users or not and then devises a factor-aware sampling strategy to sample the appropriate by comparing the relevant factors while ensuring the similarity of irrelevant factors.
3.2.5. Interpolation DNS
Interpolation DNS synthesizes the informative negative samples by injecting information from positive samples with the interpolation technology (Huang et al., 2021; Shi et al., 2022). These approaches aim to explore the balance between maintaining the informative knowledge while avoiding excessive disturbance to the distribution of the representation space during optimization, allowing recommender to adapt to the changing requirements and maintain the balance between positive and negative samples. MixGCF (Huang et al., 2021) interpolates different proportions of positive embeddings into item candidates, optimizing the representation of the selected samples and translating the potential FNS into negative samples to avoid training bias. SDNS (Shi et al., 2022) proves that the positive-dominated mixing in MixGCF has the same effect as scaling the sigmoid function in BPR loss function in DNS, and specifically proposes an improvement scheme to define a soft BPR loss function, which not only preserves the integrity of the DNS (Zhang et al., 2013), but also yields remarkable improvements in terms of effectiveness and robustness.
3.2.6. Mixed DNS
Mixed DNS combines multiple DNS strategies to generate more diverse and effective negative sample candidates (Chen et al., 2017; Lei et al., 2020; Hidasi and Karatzoglou, 2018; Hidasi et al., 2015; Huang et al., 2020; Yang et al., 2020c; Xu et al., 2021). By integrating the uniform negative examples with those above DNS strategies, the existing Mixed DNS strategy provides a flexible framework that allows recommenders to customize the task-specific combination of different DNS strategies and attempt to achieve a more comprehensive and robust negative sampling process. As the pioneering Mixed DNS strategy, NNCF (Chen et al., 2017) proposes three novel sampling strategies to significantly improve training efficiency as well as the recommendation performance, named stratified sampling, negative sharing, and stratified sampling with negative sharing. Stratified sampling selects stratum (set of links on the bipartite graph sharing the same source or destination node) based on either item or user. Negative sharing constructs negative pairs by connecting non-linked users and items in the batch without any increasing computational burdens. And stratified sampling with negative sharing denotes the combination of the above two sampling strategies. For Conversational Recommendation, EAR (Lei et al., 2020) designs a mixed negative sampling strategy to simultaneously learn the user’s general and specific preference with two types of negative samples (the traditional randomly selected samples and the samples satisfying the partially known preference in the conversation). FB-MNS (Huang et al., 2020) discovers that models trained with hard negative samples cannot outperform random negative samples. Therefore, It proposes a mixed sampling strategy to blend random and hard negatives with the ratio as in the training process, where the hard negative samples are selected between the rank position 101-500 to achieve the best Recall metric. GRU4Rec (Hidasi and Karatzoglou, 2018) extends the in-batch negative sampling strategy of GRU4Rec (Hidasi et al., 2015) with the additional shared samples in the mini-batch, where the shared samples are sampled with a popularity-based SNS strategy. MNS (Yang et al., 2020c) employs the same mixed sampling strategy as GRU4Rec (Hidasi and Karatzoglou, 2018) that incorporates the randomly selected additional samples as the supplement of the in-batch negative sampling strategy to tackle the selection bias and the inflexibility to adjust sampling distribution. Taking into account user feedback on attributes, FPAN (Xu et al., 2021) employs DNS (Zhang et al., 2013) to select informative negative samples in addition to directly sampling the non-interacted items as negative items, accelerating the training process.
| Algorithms | Sampling criteria | Advantages | Challenges | |||||||
| Universal DNS (Wu et al., 2019b; Yang et al., 2020a; Shi et al., 2023; Brin and Page, 1998; Ma et al., 2023) |
Accuracy | |||||||
| User-similarity DNS (Wu et al., 2019b; Wang et al., 2021b; Giobergia, 2022; Chen et al., 2019b) |
| Knowledge-aware DNS (Lian et al., 2020b; Chen et al., 2023b; Wang et al., 2021d) |
| Distribution-based DNS (Zhu et al., 2022; Ding et al., 2020; Lai et al., 2023) |
| Interpolation DNS (Huang et al., 2021; Shi et al., 2022) |
| Mixed DNS (Chen et al., 2017; Lei et al., 2020; Hidasi and Karatzoglou, 2018; Hidasi et al., 2015; Huang et al., 2020; Yang et al., 2020c; Xu et al., 2021) |
Flexible |
Summary: The comprehensive examination of the emphasis, advantages, and challenges of six types of DNS algorithms is detailed in Table 3. Among these methodologies, Universal DNS stands out as an easy-to-deploy and universal method. However, its reliance on user-item matching scores poses a challenge to accuracy. User-similarity DNS prioritizes identifying user behavior similarities, while facing challenges in user association dependency, particularly for new users. Knowledge-aware DNS emphasizes samples with similar attributes to positives to capture their content correlations, introducing an additional dependency on knowledge. Distribution-based DNS analyzes the inherent distribution in the dataset, focusing on real negatives but adding the dependency on space complexity. Interpolation DNS balances positive and negative samples by injecting information from positives but faces the over-smoothing issue, neglecting crucial samples. Mixed DNS has exhibited significant flexibility by combining multiple strategies, but the complex hyperparameter tuning and substantial computational costs bring new challenges to these strategies.
3.3. Adversarial Negative Generation
Adversarial Negative Generation (ANG) is a sophisticated technique which devise to enhance the robustness and performance of RS. Same as other negative sampling strategies, the core of ANG seeks to address the imbalanced data issue, where positive interactions tend to dominate while real negative instances are relatively sparse. Existing ANG studies can be fundamentally categorized into two distinct paradigms: Generative ANG and Sampled ANG. In the forthcoming sub-sections, we will expound upon these two methods from the existing ANG studies in RS.
3.3.1. Generative ANG
Generative ANG attempts to capture the underlying data distribution and generate the mendacious but plausible samples with a generative model instead of imposing any assumption on existing data. It can maintain the Nash equilibrium between generator and discriminator in generative adversarial network (GAN) (Goodfellow et al., 2014) to generate negative samples that are similar to positive samples but have subtle differences, making them hard to distinguish from positive samples (Wang et al., 2018b; Gao et al., 2019; Sun et al., 2020; Ding et al., 2019; Hao et al., 2021a; Xie et al., 2022b; Chen et al., 2022b; Zhou et al., 2021). Prevailing research in Generative ANS generally employs three methods for generating negative samples: user-specific generation, distribution-associated generation, and content-aware generation. Subsequently, we will elucidate each of them in accordance with relevant works.
User-specific generation. Generative ANG leverages the adversarial mechanism to generate the targeted negative samples for different users by generators and maximize the objective function between real and generated item representations by discriminators, thereby pushing the recommender to its performance limit. RNS (Ding et al., 2019) designs a reinforced negative sampler to generate exposure-alike negative instances and train a better recommender as the predictor through the feature matching technique and the adversarial training paradigm instead of directly choosing from exposure data. Different from RNS, ELECRec (Chen et al., 2022b) regards recommendation task as a discriminative task rather than a generative task, that is, training the discriminator as the final recommender to distinguish if a sampled item is a ‘real’ target item or not. It defines the generator as an auxiliary model of the next-item prediction task to generative high-quality plausible samples, improving the discrimination ability of the discriminator from the intricate correlations.
Distribution-associated generation. Generative ANS also adopts the adversarial training paradigm to model the optimal data distribution from the correlation between users and items. For example, DRCGR (Gao et al., 2019) generates the relevancy distribution of targeted negative samples for different users with the generator and achieves a better estimation of the relevancy between the generated negative feedback sequence and the relevant positive feedback with the discriminator, thereby automatically learning the optimal recommendation strategies. To learn the true relevance distribution of users and items, PURE (Zhou et al., 2021) continuously samples on both users and items in the representation space. Specifically, given a positive instance , the generator of PURE generates a fake item that is highly likely to be relevant to and a fake user that is likely to be relevant to from the random Gaussian noise.
Content-aware generation. With the contextual information, generative ANG can generate negative samples that align more closely with the content of positive instances, drawing from the heterogeneous knowledge encapsulated in positive samples and historical user interactions. For new-item recommendation, LARA (Sun et al., 2020) designs a generative model with multi-generators to generate the user profile which is not only close to the real profile but also relevant to the given item with the item’s attributes. AFE (Xie et al., 2022b) defines the generator to take the common features (e.g., user historical behaviors, item features, and contexts) rather than future features for generating “fake” items which may be clicked to confuse the future-aware discriminator, thus preventing the feature inconsistency and overfitting issues.
3.3.2. Sampled ANG
Sampled ANG tries to fit the underlying distribution as much as possible, and sample the most likely items as negative instances from the captured distribution (Wang et al., 2018b; Hao et al., 2021a; Wang et al., 2017b, 2018a). Different from the typical Generative ANG methods, the generator of Sampled ANG works directly select instances from the existing items in the corpus rather than generating item representations. As one of the most famous Sampled ANG works, IRGAN (Wang et al., 2017b) stands as a pioneering endeavor that incorporates GANs into RS with a game theoretical minimax game which optimizes both generator and discriminator iteratively. The primary objective of its generator lies in strategically selecting negative instances from the candidate pool and intermixing them with positive examples to confound the discriminator. Conversely, its discriminator is designed to distinguish positive samples from the negative instances sampled by the generator, and subsequently provide constructive feedback to refine the generator. To be specific, IRGAN tries to sample negative instances based on the estimated correlation between candidate items and queries vie its generator. Similarly, the generator in GraphGAN (Wang et al., 2018a) attempts to approximate the underlying connectivity distribution in the graph structure and samples the relevant vertices similar to ground truth’s real immediate neighbors to deceive the discriminator. In contrast, its discriminator tries to detect whether the sampled vertex is from ground truth or selected by the generator. NMRN-GAN (Wang et al., 2018b) also develops a GAN-based noise sampler to adaptively sample the hardest negatives from the randomly selected candidates, which considers both the specific user and the recommender parameters with the policy gradient-based reinforcement learning framework. AFT (Hao et al., 2021a) proposes a domain-specific masked encoder and a GAN-based two-step feature translation to aggregate the domain-specific preferences from other domains, thereby retrieving the fake clicked samples in different domains for adversarial training.
| Algorithms |
Advantages | Challenges | |||||||
| Generative ANG (Gao et al., 2019; Sun et al., 2020; Ding et al., 2019; Xie et al., 2022b; Chen et al., 2022b; Zhou et al., 2021) |
| Sampled ANG (Wang et al., 2018b; Hao et al., 2021a; Wang et al., 2017b, 2018a) |
Summary: We systematically outline the generation strategies, advantages, and challenges of each ANG strategy in Table 4. Here, Generative ANG endeavors to leverage GAN to generate the mendacious yet plausible samples, in which the generator and discriminator play a game-theoretical minimax game. The generator attempts to synthesize the plausible negative instances, while the discriminator seeks to distinguish between the genuine and generated samples. Through the iterative adversarial training, the generator becomes increasingly adept at generating mendacious but plausible negative instances, thereby enabling diverse negative generation, improving recommender generalization. However, it faces the risk of pattern breakdown and may lead to the relatively complex training process. In contrast, Sampled ANG strategically samples appropriate items from the corpus with the Nash equilibrium state between the generator and discriminator. Its advantage lies in guaranteeing the simplified GAN training procedure and the utilization of instances originating exclusively from authentic data. This eliminates any potential training bias stemming from varying levels of GAN’s generation complexity. Consequently, it effectively alleviates the imbalance between positive and negative samples in RS, enabling its training on a more balanced dataset. Nevertheless, it encounters challenges such as the homogenization of negatives and the difficulty in covering all potential user preferences.
3.4. Importance Re-weighting
Importance Re-weighting (IRW) is one of the sophisticated statistical techniques employed in Data Mining (Chen et al., 2021a; Krishnan et al., 2020; Yu et al., 2020; Wu and Ye, 2020). Its core revolves around the nuanced adjustment of sample weights to prioritize the significance of certain important negative samples (Chen et al., 2023d; Zhuo et al., 2022). For the recommendation task where unobserved samples vastly outnumber positive feedback, IRW can be instrumental in rebalancing the real-world dataset by assigning diverse weights to different samples, thus enabling the algorithm to focus its learning efforts on the rare but crucial negative instances (Wang et al., 2021b; Chen et al., 2023b). Its essential idea is consistent with the memorization effect: recommenders tend to learn accessible user preferences in the early stage of the training process and eventually memorize all complicated interactions. Properly assigning weights to individual samples attunes recommenders to the intricacies of imbalanced implicit datasets, thus recalibrating the learning process for the disparate data distribution. This adaptive weighting not only addresses the challenge of imbalanced data but also fosters the development of more robust and precise recommenders, ultimately bring in superior prediction accuracy. The investigation of IRW strategies classifies them into Attention-based IRW, Knowledge-aware IRW, and Debiased IRW and we will introduce their technical details in the subsequent sub-sections.
3.4.1. Attention-based IRW
Attention-based IRW assigns importance scores and adjusts weights for each item, which could focus on the most relevant items for each user and provide suitable recommendations aligned with the user’s personalized interests with the attention mechanism. (Gao et al., 2022; Chen et al., 2021a; Dai et al., 2024b; Chen et al., 2023d; Zhuo et al., 2022). To efficiently train recommenders with millions of items in the real-world applications, this strategy attempts to sample negative candidates from the uniform distribution and then correct them via importance weighting. A series of works attempt to focus on the safe (easy) negative samples while overlooking the hard negative samples. To adaptively prune the large-loss interactions, ADT (Wang et al., 2021c) devises the truncated loss paradigm to discard the large-loss samples (hard ones) with a dynamic threshold and a reweighted loss paradigm to adaptively lower the weights of these samples. UIB (Zhuo et al., 2022) innovatively introduces an auxiliary score for each user to define a personalized decision boundary and individually penalize samples that cross the boundary with a hybrid loss of point-wise and pair-wise paradigm. Moreover, some studies assume that more informative negative samples can contribute more to the gradient, the magnitude of the gradient becomes much larger so that training can be accelerated. To assist the learning process from the most important historical transitions, DEERS (Zhao et al., 2018) proposes a prioritized sampling strategy to greedily select the most informative negative sample while ensuring that the probability of being sampled is monotonic in a transition’s priority in reinforcement learning.
Recently, some studies imitate the learning attention in human perception to assign different weights to diverse negative samples from simple to complex (whose informativeness is from less to more). PRIS (Lian et al., 2020a) proposes a personalized negative sampler based on rejection sampling to assign larger weights to more informative negative samples, which can contribute more to the gradient of the ranking loss function and accelerate the convergence rate. To address the distribution mismatch, REINFORCE (Chen et al., 2019a) takes the first-order approximation and ignores the state visitation differences as the importance weights of diverse negative samples with the normalized importance sampling strategy to construct a slightly biased estimator of the policy gradient with lower variance. SGDL (Gao et al., 2022) investigates the memorization effect within several representative recommenders’ learning process, delineates them into the noise-resistant period and the noise-sensitive period. Then it collects memorized interactions as denoising signals at the noise-resistant period and imposes a weight on each sample loss to enhance the robustness of the noise-sensitive period. GeoSAN (Lian et al., 2020b) also proposes a weighted BCE loss based on importance sampling to reweight the unvisited locations with the negative probability, enabling more informative locations can be more emphasized even with the uniform negative sampler. ContraRec (Wang et al., 2023a) leverages the temperature hyper-parameter in the proposed CTC loss to assign diverse attention to negative samples according to their hardness in prediction.
3.4.2. Knowledge-aware IRW
Attention-based IRW studies typically design a reweighting function that maps each sample’s score or loss value to its learning weight, thereby enabling the recommender to adaptively learn the information of each negative sample in model optimization (Gao et al., 2022; Wu et al., 2022a). However, Knowledge-aware IRW argues that users’ social contexts and items’ heterogeneous information serve as influential determinants in shaping the reciprocal engagements between users and items (Yang et al., 2023b; Kim et al., 2021a; Krishnan et al., 2020; Chen et al., 2019b). In tandem, these elements covertly impact the confidence of the data. When an item is more popular among this user’s social neighbors, the user is more likely to know the item, and his feedback is more attributed to his real preference. Correspondingly, in the situation that user’s historical behaviors exhibit a level of coherence in their content, recommender can effortlessly predict the items sharing the similar semantic attributes (Chen et al., 2021a; Wang et al., 2021b). This, in turn, addresses the user’s preference for particular content within the contextual paradigm, and this knowledge will be more reliable in deriving user preference.
Within the realm of recommendation, the pre-defined knowledge is typically categorized into user-oriented social knowledge and item-oriented heterogeneous knowledge. Different knowledge-aware IRW methods are commonly utilized to integrate these distinct forms of knowledge into importance re-weighting. For the user-oriented social knowledge, SamWalker (Chen et al., 2019b) simulates item information propagation along the real social network and adaptively specifies individual confidence weights to different interactions based on user’s social contexts. To mitigate the reliance on authentic user affiliations, SamWalker++ (Wang et al., 2021b) introduces the rationale of “wisdom of the crowds” to define the pseudo-social network, in which similar users are connected with specific item nodes or community nodes. Under this scenario, SamWalker++ can model data confidence and assign adaptive confidence weight to each interaction without any side information. Furthermore, for the item-oriented heterogeneous knowledge, PREMERE (Kim et al., 2021a) proposes a meta-learning-based reweighting strategy to produce the most suitable weight at each optimization state, which can handle the rarity of positive samples and the noisiness of negative samples simultaneously. URL (Chen et al., 2021a) leverages an auxiliary task to consider user state and context knowledge in reinforcement learning and reweighs each item with a long-term reward to avoid the high variance in policy learning caused by the inverse-propensity weighting. MMT-Net (Krishnan et al., 2020) propose an inverse novelty measure, noted as context-bias, to attenuate the original supervised loss by emulating variance-reduction, thereby inducing a novelty-weighted strategy focused on harder samples as training proceeds.
3.4.3. Debiased IRW
Debiased IRW identifies and corrects the ubiquitous biases inherent in RS (e.g., popularity bias, exposure bias) and assigns higher weights to items that have been overlooked in the past to deliver more equitable and diverse recommendations (Shen et al., 2021; Yu et al., 2020; Chen et al., 2023b). It operates on the premise that traditional recommenders may inadvertently exacerbate disparities by assigning undue weight to certain categories or user preferences.
By assigning diverse weights to items with various categories, debiased IRW endeavors to counteract any predispositions within traditional recommenders. It can foster more equitable item representation, ensuring that no certain categories disproportionately influence the prediction results. For example, FairNeg (Chen et al., 2023b) proposes a negative sampling distribution mixup mechanism, which incorporates both importance-aware and fairness-aware sampling distribution to simultaneously strengthen the item representation and item-oriented group fairness. Furthermore, some debiased IRW works also meticulously design and incorporate a suite of fairness metrics and concepts to evaluate the fairness and mitigate the inherent bias. SAR-Net (Shen et al., 2021) designs a concept of “Fairness Coefficient” to measure each sample’s importance and use it to reweigh the prediction of the debiased expert networks, thereby mitigating the data fairness issue caused by manual intervention. IF4URec (Yu et al., 2020) introduces the “Influence Function” to measure the effect on the estimator when adding a small perturbation on training sample, which could reveal its importance. It is able to utilize “Influence Function” to reweigh the training loss of each sample to get less loss in an unbiased validation for debiasing.
| Algorithms | Reweighting functions | Advantages | Challenges | |||||||
| Attention-based IRW (Lian et al., 2020a; Wang et al., 2021c; Zhao et al., 2018; Gao et al., 2022; Chen et al., 2021a; Covington et al., 2016; Wu et al., 2022a; Chen et al., 2023d; Zhuo et al., 2022) |
| Knowledge-aware IRW (Kim et al., 2021a; Krishnan et al., 2020; Chen et al., 2021a; Wang et al., 2021b; Chen et al., 2019b) |
| Debiased IRW (Shen et al., 2021; Yu et al., 2020; Chen et al., 2023b) |
Summary: We have systematically delineated the reweighting functions, advantages and challenges of these above IRW strategies in Table 5. Here, Attention-based IRW emphasizes assigning diverse importance to each item with user interest attention, making it adaptably and capable of flexibly adjusting weights. However, it faces challenges due to its complexity and relatively poor interpretability. Knowledge-aware IRW focuses on identifying each item’s importance with external structured knowledge, offering suitability for cold-start issues but introducing the additional dependency on knowledge. Debiased IRW aims to correct ubiquitous biases, assigning higher weights to overlooked items to provide equitable and diverse recommendations. Nevertheless, it involves the trade-off between fairness and precision.
3.5. Knowledge-enhanced Negative Sampling
Knowledge-enhanced Negative Sampling (KNS) is one of the pivotal negative sampling strategies in recommendation, which leverages the availability of supplementary informative knowledge, such as the auxiliary information (i.e., users’ social contexts and items’ heterogeneous knowledge) and knowledge graphs, to refine the selection of negative samples. It operates within a broad spectrum of recommendation scenarios, ranging from context-aware recommendation to graph-enhanced recommendation, where knowledge plays a crucial role. The core of Knowledge-enhanced Negative Sampling unfolds in two distinctive dimensions: General KNS and KG-based KNS.
3.5.1. General KNS
The rich auxiliary relationships of users and the informative knowledge of items typically imply user preferences and item semantics, which could help overcome the sparsity issue in RS. General KNS selects negative samples based on the existing latent knowledge patterns to better understand the semantics and relations between users and items (Steck, 2018; Sun et al., 2015; Zhao et al., 2019; Krishnan et al., 2020). The learning process of the knowledge-based recommendation is grounded on contextual information. For instance, MMT-Net (Krishnan et al., 2020) designs the context-bias to reveal the novel tuples which display uncommon combinations of the interaction and its corresponding context, which can be defined as , where denotes the final context combinations. To enhance the training effectiveness of the proposed context-bias, MMT-Net develops a negative sampling approximation with two learning objectives to randomly identify the likely item given the user and interaction context and identify the likely context given the user and the item, respectively. Then the context-bias has been used to attenuate the loss as , thereby inducing a novelty-weighted training curriculum focused on harder samples as training proceeds. SA_OCCF (Sun et al., 2015) leverages sentiment information extracted from user reviews to assign the weight for each negative example rather than the global weighting scheme. Instead of the review information, some works also leverage the informative genre and category knowledge for negative samplings. To list a few, LDA (Steck, 2018) devises a two-tiered sampling approach that first samples a genre from the learned genre-distribution for user , and then samples the movie as the negative sample from the learned distribution regarding the genre . IntentGC (Zhao et al., 2019) samples negative items in the same leaf category as the corresponding positive item to make the model capable of distinguishing hard items.
3.5.2. KG-based KNS
KG-based KNS leverages the explicit association and high-order correlation among users, items, and other entities in KG to sample negative instances (Liu et al., 2021b; Ali et al., 2020; Wang et al., 2020; Chen et al., 2022a). KG, which introduces extra relations among items and real-world entities (e.g., item attributes) with the comprehensive structural knowledge, are well known for its potential to enhance recommenders’ accuracy and explainability. To precisely improve the semantic relevance between positive and negative samples and enhance the accessible information of the sampled negative instances, a series of works attempt to introduce the high-order structural coherence and the multi-hop node connectivity into negative sampling in RS (Liu et al., 2021b; Ali et al., 2020; Wang et al., 2020). For example, AnchorKG (Liu et al., 2021b) design a simple but effective strategy to randomly select the in and conduct a random walk along a -hop path starting from the positive sample in KG. To minimize computational complexity, PR-HNE (Ali et al., 2020) employs unigram distribution to sample items (nodes) in candidates that appear in the KG independently of the presence of other items over each interaction (edge). With the reinforcement learning (RL) paradigm, KGPolicy (Wang et al., 2020) samples knowledge-enhanced negative items from the relationships between items and entities in KG, emphasizing KG entities and items linked within two-hop paths and facilitating recommenders’ learning process. KGAttack (Chen et al., 2022a) proposes knowledge-enhanced candidate selection to employ KG for localizing some relevant item candidates with similar attributes. Subsequently, it leverages RL paradigm to obtain the sampling probabilities of each candidate item by calculating the similarity between the current state representations and graph embeddings.
| Algorithms | Knowledge source | Advantages | Challenges | |||||||||
| General KNS (Steck, 2018; Sun et al., 2015; Zhao et al., 2019; Krishnan et al., 2020) |
| KG-based KNS (Liu et al., 2021b; Ali et al., 2020; Wang et al., 2020; Chen et al., 2022a) |
Summary: As demonstrated in Table 6, General KNS exploits all available heterogeneous knowledge to sample the more informative negative samples with richer semantic meaning, thereby outperforming the traditional negative sampling methods. It ensures that the selected negative samples are not only dissimilar to the positive samples but also incorporate meaningful semantic relationships with the positive samples, aligning more closely with users’ preferences. However, it relies on high-quality knowledge and additionally introduces data processing complexity. In contrast, KG-based KNS leverages the structured information from KG to enhance the quality of the sampled negatives. By incorporating the user-item interactions to the entities in KG, KG-based KNS can introduce the high-order structural coherence and the multi-hop node connectivity into negative samples, thereby precisely improving the semantic relevance rather than the feature relevance between positive and negative samples. However, it excessively depends on KG construction and faces challenges related to the computational complexity of modeling and maintaining the KG.
4. Negative Sampling in Multiple Practical Recommendation Scenarios
The field of recommender systems is replete with a vast spectrum of diverse scenarios, which stems from user demands, historical behaviors, and heterogeneous applications, thereby incorporating the field with a blend of challenges and prospects (Gao et al., 2024; Chen et al., 2023a). Benefits from its superior performance and easy-to-employ character, negative sampling has garnered significant attention from researchers and practitioners in recommendation (Zhang et al., 2013; Shi et al., 2023; Ding et al., 2019; Lian et al., 2020a). Some researchers have attempted to design a universal negative sampling method that is suitable for all recommendation scenarios (Ma et al., 2023). Nonetheless, the employment, objective, and data availability significantly vary across diverse scenarios, leading to the absence of a negative sampling algorithm that can comprehensively encompass all recommendation scenarios. In this section, we undertake an in-depth exploration of the distinctiveness of various recommendation scenarios, delineating subtle disparities in the negative sampling methodologies under the same scenarios. Ultimately, we conclude with a comprehensive discussion and outlook concerning the paradigm of tailored negative sampling techniques for each specific scenario. Specifically, the subsequent discussions span six pivotal recommendation scenarios in the following subsections: Collaborative-guided Recommendation, Sequential Recommendation, Multi-modal Recommendation, Multi-behavior Recommendation, Cross-domain Recommendation and CL-enhanced Recommendation. These discrete discussions and analyses illustrate the adaptability and customization inherent to negative sampling strategies, a fundamental prerequisite for optimizing recommendation algorithms in diverse recommendation contexts.
| Recommendation Tasks | Algorithms | Negative Sampling Strategies | Venue | Year |
| Collaborative Filtering | NCF (He et al., 2017) | Uniform SNS | WWW | 2017 |
| ENMF (Chen et al., 2020c) | Non-sampling SNS | TOIS | 2020 | |
| AHNS (Lai et al., 2024) | Universal DNS | AAAI | 2024 | |
| TIL (Wu et al., 2022a) | Attention-based IRW | CIKM | 2022 | |
| DNS* (Shi et al., 2023) | Universal DNS | WWW | 2023 | |
| Graph-based Recommendation | DRCGR (Gao et al., 2019) | Generative ANG | ICDM | 2019 |
| IntentGC (Zhao et al., 2019) | General KNS | KDD | 2019 | |
| JNSKR (Li et al., 2021) | Non-sampling SNS | SIGIR | 2020 | |
| AnchorKG (Liu et al., 2021b) | KG-based KNS | KDD | 2021 | |
| SGCF (He et al., 2023a) | Uniform SNS | WSDM | 2023 | |
| Sequential Recommendation | DEEMS (Wu et al., 2019b) | User-similarity DNS | KDD | 2019 |
| GeoSAN (Lian et al., 2020b) | Knowledge-aware DNS | KDD | 2020 | |
| ELECRec (Chen et al., 2022b) | Generative ANG | SIGIR | 2022 | |
| PDRec (Ma et al., 2024a) | Universal DNS | AAAI | 2024 | |
| ContraRec (Wang et al., 2023a) | Attention-based IRW | TOIS | 2024 | |
| Multi-modal Recommendation | RPRM (Wang et al., 2021d) | Knowledge-aware DNS | WWW | 2021 |
| KGPL (Togashi et al., 2021) | Popularity-based SNS | WSDM | 2021 | |
| POWERec (Dong et al., 2024) | General KNS | Information Fusion | 2023 | |
| FairNeg (Chen et al., 2023b) | Debiased IRW | WWW | 2023 | |
| FREEDOM (Zhou and Shen, 2023) | Uniform SNS | MM | 2023 | |
| Multi-behavior Recommendation | DRN (Zheng et al., 2018) | Predefined SNS | WWW | 2018 |
| ATRank (Zhou et al., 2018) | Uniform SNS | AAAI | 2018 | |
| KHGT (Xia et al., 2021) | KG-based KNS | AAAI | 2021 | |
| MBHT (Yang et al., 2022b) | Uniform SNS | KDD | 2022 | |
| MBRec (Xia et al., 2022) | General KNS | TNNLS | 2022 | |
| Cross-domain Recommendation | MMT-Net (Krishnan et al., 2020) | Knowledge-aware IRW | SIGIR | 2020 |
| ML-MGC (Sun et al., 2021) | Predefined SNS | IJCNN | 2021 | |
| DDGHM (Zheng et al., 2022) | Scenario-aware NACL | MM | 2022 | |
| C2DSR (Cao et al., 2022) | Uniform SNS | CIKM | 2022 | |
| RealHNS (Ma et al., 2023) | Universal DNS | Recsys | 2023 |
4.1. Collaborative-guided Recommendation
Collaborative-guided Recommendation (CR) aims to gather and analyze all the collaborative behaviors for predicting users’ potential preferences (Rendle et al., 2009). It typically focuses on projecting users and items into the latent embedding space to reflect their correlation (He et al., 2017). In contrast to other recommendation scenarios, the available data of CR is exclusively restricted in the user-item interactions, which constrains its practical accuracy. Moreover, it is widely recognized that merely considering positive instances is insufficient to capture the intricacies of user preferences (Zhang et al., 2013; Shi et al., 2023). In light of this, a series of CF methods have investigated how to maximize the utilization of interaction data to sample more informative negative instances (Chen et al., 2020c; Shi et al., 2023; Ding et al., 2020; Wu et al., 2022a).
Constrained by the data availability inherent to CR tasks, prevailing negative sampling methods typically employ SNS approaches to facilitate the process of negative sample selection. Precisely, randomly sampling items from the unobserved corpus with the same probability (He et al., 2023a, 2017) are extensively employed in CR. A substantial number of CR approaches also focus on the dynamic selection of negative samples, which are information-rich during the training of CR models. DNS (Zhang et al., 2013) and DNS (Shi et al., 2023) regard the items with the highest score as the most informative samples and attempt to select them as the negative samples. SRNS (Ding et al., 2020), GDNS (Zhu et al., 2022), and DENS (Lai et al., 2023) leverage the inherent distribution (e.g., statistical information and disentangled representation) between positive and negative samples to dynamically select the suitable negative instances under the current context in training. Furthermore, a CL-based CR algorithm, EGLN (Yang et al., 2021d), attempts to conduct the structural augmentation on the interaction graph to maximize the local-global consistency in collaborative representation learning.
4.2. Sequential Recommendation
Sequential Recommendation (SR) is a straightforward but effective approach for capturing users’ dynamic preference patterns (Dallmann et al., 2021; Wang et al., 2023a; Ma et al., 2024c). Its goal is to predict the next items the user is likely to interacted with by modeling the temporal correlations in his/her behavioral sequence. Some studies have proven the effectiveness of hard negative sampling strategies in SR (Wu et al., 2019b; Chen et al., 2022b; Xie et al., 2022a). Through the selection of negative examples that exhibit a certain similarity to positive samples, these strategies bolster the recommender’s capability to finely distinguish users’ dynamic preferences. Most existing SR methods randomly select unobserved items from the corpus as the negative samples (Yang et al., 2022b; Dallmann et al., 2021; Petrov and Macdonald, 2022). With the development of contrastive learning (CL), some studies rely on diverse sequence augmentation methods (e.g., crop, mask, reorder, substitute, and insert) to generate negative samples, assisting in modeling users’ temporal preferences. Despite the utilization of hard negative sampling strategies, DEEMS (Wu et al., 2019b) and GeoSAN (Lian et al., 2020b) opt to treat each interaction with equitable importance rather than explicitly considering the temporal relationships among behaviors Similarly, ELECRec (Chen et al., 2022b) leverages discriminator in GAN to distinguish the target sample, overlooking the sequential knowledge within behavioral sequences. Evidently, negative sampling holds particular significance for SR, which has been momentarily disregarded by existing researchers and may serve as a promising avenue of exploration for future scholars.
4.3. Multi-modal Recommendation
Multi-modal Recommendation (MMR) is crucial in real-world recommendation applications with massive modality-specific information, that is, micro-videos, images, audio, and text (Wang et al., 2023a; Tao et al., 2022; Ma et al., 2024d). Since the multi-modal information may reflect users’ actual preferences within the fine-grained modality level (Chen et al., 2023b), the idealized multi-modal recommender possesses the capacity to uncover the hidden relationships between different modalities and recover the complementary information that can not be captured by CR methods (Dong et al., 2024). To achieve satisfactory performance, MMR methods typically require a substantial amount of both positive (present in the observed interactions) and negative items (obscured within the unobserved interactions) to train recommenders in a supervised manner. This motivates researchers to explore how to sample more informative negative instances that better align with users’ actual preferences with the multi-modal knowledge. Similar to other recommendation scenarios, most related studies commonly apply SNS (Wang et al., 2023a; Wu et al., 2019a; Wang et al., 2019) to select negative samples in MMR. Some works also investigate the hidden correlations within the items’ multiple modalities to conduct the modality-aware CL pairs. For the traditional negative sampling strategies, FairNeg (Chen et al., 2023b) achieves debiased recommendation via capturing the group-level unfairness and balancing the sampling probabilities of each semantic group. RPRM (Wang et al., 2021d) attempts to model the review properties’ consistency among different items and consequently sample the negatives with similar review properties of positive samples. POWERec (Dong et al., 2024) leverages the strength of multi-modal representations learned from the identical item to enhance user interest understanding in the unreliable modality, making all the modality-specific user interests to be fully learned. These algorithms can incorporate the multi-modal information to intelligently refine the selection of the relevant negative instances, which exhibits meaningful semantic relationships with the positive samples and closely aligns with the user’s intricate interests.
denotes “like”,
denotes “dislike”,
denotes “add to cart”,
denotes “purchase”,
denotes “observed real behaviors” and
denotes “unobserved samples”.
4.4. Multi-behavior Recommendation
As illustrated in Fig. 8, real-world RS typically involves the concept of real negative feedback in addition to implicit and explicit feedback, which contributes to holistic understanding of user preferences (Xie et al., 2021; Huang et al., 2023). It refers to the authentic user negative behaviors, with its definition intimately linked to various recommendation scenarios. For instance, non-purchases are considered negative samples in the context of purchases, while in the context of clicks, exposures without clicks constitute negative samples. Similarly, non-likes, in comparison to likes, are treated as negative samples. This encapsulates explicit information of users’ disapproval, wielding significant influence in recommendation (Ding et al., 2023).
Benefits from the multiple types of interaction between users and items in real-world scenarios, multi-behavior recommendation (MBR) refers to taking full advantage of these heterogeneous user behaviors (e.g., skipped, clicked, and ordered) to alleviate the data sparsity or cold-start issues (Xie et al., 2021; Xia et al., 2022; Gu et al., 2021; Zhang et al., 2024a). As illustrated in Fig. 9, its core idea is to model the cross-type behavior dependency in RS, which can provide complementary information for user’s interest encoding (Gu et al., 2021; Zhou et al., 2018; Yang et al., 2021a). In such an application, effectively modeling the multi-type user behaviors can provide auxiliary knowledge and precise user preference to characterize the underlying semantics of user’s actual interest representation (Zheng et al., 2018; Gu et al., 2022). Existing MBR works primarily consider the multi-behavior dependencies with the predefined correlations (Yang et al., 2022b; Xia et al., 2022), while seldom embarking on the selection of diverse negative samples for individual behaviors through negative sampling. To some extent, this limits its ability to capture the intricate cross-type behavior dependencies within complex real-world RS scenarios. For instance, ZEUS (Gu et al., 2021) and HMG-CR (Yang et al., 2021a) independently construct negative CL pairs by forming inter-connections between multiple behaviors and generating behavior-aware hyper meta-graphs. It enables the simultaneous modeling of the diverse interaction patterns and underlying cross-type behavior inter-dependencies, verifying the positive effects of incorporating the cross-type of multi-behavioral context in MBR.
4.5. Cross-domain Recommendation
Cross-domain recommendation (CDR) is one of the representative methods to alleviate the data sparsity problem and the cold-start problems in traditional RS with the auxiliary user behaviors collected from other domains (Cao et al., 2022; Ma et al., 2024b). It aims to leverage the correlation of users’ multi-domain behaviors to accurately transfer the informative knowledge from the source domains and model users’ personalized preference in the target domain (Bi et al., 2020). Traditional CDR methods typically conduct the cross-domain knowledge transfer by modeling the feature-level cross-domain correlations with aligned constraint (Krishnan et al., 2020; Liu et al., 2020), adversarial training (Liu et al., 2021a) and contrastive learning (Zheng et al., 2022; Ma et al., 2024b). However, these methods merely focus on the randomly-selected negative samples from the target domain, ignoring the cross-domain differences at the sample level. This may disregard users’ preference discrepancy between multiple domains and leads to the sub-optimal performance. DDGHM (Zheng et al., 2022) utilizes random mask operator for negative pair augmentation in CL optimization. Nevertheless, when it comes to recommendation tasks, the approach of random sampling within the target domain is still employed. RealHNS (Ma et al., 2023) is the pioneer to systematically discover the false and refine the real from all HNS in CDR. It conducts the multi-grained HNS selectors to find high-quality general and cross-domain HNS with a curriculum learning framework to alleviate cross-domain negative transfer. Notably, there exists substantial uncharted territories for negative sampling methods in CDR, which awaits the investigation by future researchers. This encompasses various areas, including, but not limited to, KNS, ANG, and Debiased IRW.
4.6. CL-enhanced Recommendation
Inspired by the success of CL in computer vision (Han et al., 2021), natural language processing (Kim et al., 2021b) and graph representation learning (Mo et al., 2022), some works (Shuai et al., 2022; Wang et al., 2022; Zhou et al., 2020; Tang et al., 2024a) apply CL paradigm in RS. CL aims to maximize the agreement between positive pairs (an item’s representation and its corresponding augmented representations) and minimize it between negative pairs (an item’s representation and other items’ augmented representations) to acquire the transferable knowledge between the original view and the augmented view. Intuitively, CL relies heavily on its negative augmentation (NA) strategy, which is due to the fact that the appropriate augmented view can facilitate the recommender to exploit richer underlying semantic correlation. The most widely used augmentation strategy in CL-based recommendation is the in-batch negative sampling. It treats the augmented representations of the remaining items in the mini-batch where the positive example is located as negative examples, rather than selecting from the corpus, thereby aiding in accelerating its training process without the imposition of additional computational complexity.
To deeply understand the negative augmentation strategies in CL-based recommendation, we comprehensively dissect its core methodology into four distinct dimensions: pre-defined NA, graph-based NA, sequence-based NA and scenario-aware NA. Firstly, pre-defined NA fundamentally aims to refine CL process by augmenting the negative sample pool based on the pre-defined criteria (Du et al., 2022; Zhang et al., 2021b; Gu et al., 2022), which are typically derived from task-specific expertise. For instance, CauseRec (Zhang et al., 2021b) models the counterfactual data distribution by replacing the identified indispensable concepts in user behavioral sequences as negative pairs. Secondly, graph-based NA refers to generating the negative pairs based on distinct types of graph structures, including interaction graphs, social graphs, and heterogeneous graphs (Zhang et al., 2021a; Wu et al., 2021a; Yang et al., 2021d; Ma et al., 2022). It typically perturbs the nodes and edges to provide a nuanced understanding of user-item interactions. SGL (Wu et al., 2021a), EGLN (Yang et al., 2021d), and RGCL (Shuai et al., 2022) fully consider the correlations between nodes and edges within the graph structure to propose several graph-based NA strategies, including node (edge) dropout, node (edge) discrimination, random walk, and structure perturbation. Thirdly, sequence-based NA focuses on augmenting negative samples (e.g., shuffle, mask, crop, etc.) derived from historical behavior sequences (Zhou et al., 2020; Tian et al., 2022; Xie et al., 2022a; Wang et al., 2022). Its effective mechanism resides in comprehensively capturing the inherent dynamic characteristics of user behavior sequences. With the tailored self-supervised representation augmentation, it effectively handles the sequence structure and builds more robust and accurate recommenders. Finally, scenario-aware NA involves tailoring the augmentation of negative samples to accommodate the intricacies and demands of various recommendation scenarios. Here, MMR derives multi-view relationships of the user’s preferences and item’s representation across multiple modalities with CL to avoid the problem of limited diversity or bias (Wei et al., 2021; Tao et al., 2022), and MBR constructs negative examples based on the auxiliary behavior data to model the different user behavior patterns through CL (Gu et al., 2021; Yang et al., 2021a). Adaptively selecting the appropriate NA strategy based on the scenario requirements and data availability in CL-based recommendation is sensible. Meanwhile, how to construct a universal augmentation method to serve various recommendation scenarios remains an urgent issue.
5. Conclusion and Future Direction
Sampling the appropriate negative instances that align with the current model optimization state to model users’ unbiased preference is the fundamental yet challenging issue in recommendation. In this survey, we conduct a comprehensive and systematic investigation on negative sampling in recommendation over the past ten years. By holistically analyzing the significance and the challenges of negative sampling research in recommendation, we collect the existing related studies, organize and cluster them into five categories, and outline their diverse sampling schemes to discuss the current research contributions within this area. Subsequently, we systematically detail the diverse insights of the tailored negative sampling approaches in different recommendation scenarios from the perspective of the specific recommendation objective. Finally, we will discuss several important yet not-well-explored research directions in the field of negative sampling to inspire some future studies. We hope this survey is able to offer a comprehensive understanding of this promising but easy-to-ignored area to both newcomers and experts from academia and industry, who are dedicated to negative sampling research in recommendation and to provide some insights for potential future research.
5.1. Further Explorations on False Negative Sample Issue
The primary objective of existing negative sampling algorithms in RS is to strategically sample informative yet not excessively challenging negative instances (RHNS) from item candidates (Shi et al., 2023; Ma et al., 2023). These negative samples are then incorporated to augment the training process of the recommender. From the former studies, we have ascertained the presence of a rather indistinct boundary between SNS and HNS within the prevailing negative sampling strategies (Yang et al., 2020a; Zhu et al., 2022; Ding et al., 2020). This implies that a degree of randomness is inherent in these approaches during negative sampling, rendering it challenging to precisely address the issue of false negatives. Furthermore, users occasionally exhibit preferences for some specific items that diverge significantly from their behavioral patterns, making it difficult to precisely model this occasional preference with user’s implicit behaviors (Guo et al., 2020a; Yin et al., 2019). Following the established nomenclature in RealHNS (Ma et al., 2023), this type of samples is defined as occasional HNS. Regarding the occasional HNS, the utilization of distinct users’ multiple behaviors and items’ heterogeneous knowledge across different recommendation scenarios may aid in modeling their occasional preference. This could effectively facilitate the utilization or denoising of occasional HNS.
Taking the classical hard negative sampling method (DNS (Zhang et al., 2013)) as an illustrative example, the size of the candidates effectively determines the “hard” level of the sampled negative samples. In fact, the concept of “hard” possesses varying definitions across diverse recommenders and even distinct datasets. The indiscriminate employment of DNS to varying recommenders and datasets has been proved infeasible, manifesting the absence of interpretability and stability (Zhou et al., 2021; Chen et al., 2021d; Li et al., 2021). Consequently, elevating the interpretability, stability, and scalability of existing hard negative sampling strategies while concurrently maintaining the precision remains an open-ended issue and deserves further exploration.
5.2. Curriculum Learning on Hard Negative Sampling
Inspired by the empirical evidence in behavior and cognitive science literature, the curriculum learning scheme has been introduced to improve the generalization capacity and convergence rate of neural networks (Peterson, 2004; Bengio et al., 2009). The core of curriculum learning is to train the model with easier candidates and then gradually increase the hard level of samples until its convergence (Wang et al., 2021a). This process imitates the learning order in human curricula. Its fundamental objective is to enhance the model’s learning efficiency by providing more informative training signals to the model as it gradually captures and refines specific knowledge (Wang et al., 2021a). With the assumption that indiscriminately introducing all HNS at the beginning of training may lead to computational wastage, sub-optimal performance, and excessively high gradient magnitudes, some recommendation algorithms attempt to select negative samples with the curriculum learning paradigm (Ma et al., 2023; Yang et al., 2021c). Curriculum learning is able to gradually incorporate increasingly hard training instances in recommender’s learning process. It aligns seamlessly with the negative sampling algorithms, which are dedicated to the selection of the most informative samples within the item candidates, those that promise the most significant model enhancement (Chen et al., 2021b). In this situation, negative sampling algorithms and curriculum learning paradigm exhibit similarity to some extent, as they both focus on directing model learning in dynamic and intricate environments (Chen et al., 2021e). Such inherent resemblances present us the opportunity to intricately amalgamate them to guide the training process of recommender more effectively, ultimately culminating in improved performance and enhanced training efficiency.
In the not-so-distant future, it is imperative to explore the integration of negative sampling algorithms with the curriculum learning paradigm. Confronted with the increasingly intricate scenarios and diverse behaviors, the exploration of refining existing negative sampling methods to bolster the seamless collaboration between them and curriculum learning becomes profoundly pivotal. Additionally, the exploration of diverse training strategies holds the potential to achieve heightened model performance and greater efficiency.
5.3. Causal Inference for Negative Sample Understanding
Causality is the science of cause and effect in which the effect partly originates from the cause to some extent. As mentioned in (Gao et al., 2024), causal inference refers to the process of determining and further leveraging the causal relation from the experimental results or observational data. In the recent past, causal inference has been introduced into RS (Yang et al., 2021b) to extract the causality within the recommenders in modeling intricate user preferences, departing from the traditional correlation-centric paradigm (He et al., 2022, 2023b). This paradigm aims to determine how user historical behaviors influence the recommendation results, allowing for a comprehensive comprehension of user preferences.
From the perspective of negative sampling, we posit that the integration of negative sampling with causal inference holds the potential to enable the selection of interpretable and traceable negative samples which can dynamically represent the most suitable instances for recommender. Two kinds of causality widely exist in RS, user-aspect, and interaction-aspect (Gao et al., 2024). Regarding the user-aspect causality, user-similarity DNS can assist in unearthing commonalities within user behaviors by incorporating causal relationships, thus simulating the driving process behind user decision-making (Wu et al., 2019b; Wang et al., 2021b; Chen et al., 2019b). In contrast, concerning interaction-aspect causality, universal DNS and distribution-based DNS are proficient at reinforcing users’ unadulterated preferences by exploring and establishing causal links between user behaviors and recommended results (Shi et al., 2023; Zhu et al., 2022; Ding et al., 2020). Under these circumstances, negative sample selection is not solely based on the item correlation but also involves the discernment of potential causal relationships among them (Chen et al., 2019c). Nevertheless, the integration of negative sampling with causal inference remains an underexplored research avenue in RS, with only a limited number of studies delving into this domain. We believe that it holds great potential to offer the opportunity to provide more accurate, interpretable, and user-centric recommendations.
5.4. Alleviating Biases in Negative Sampling
The user behavioral data utilized in the training of recommenders originates from the observed data within the real-world RS rather than experimental data (Chen et al., 2023a). Consequently, this inherent source introduces various biases, which are driven by the diverse exposure mechanisms employed within the real-world large-scale RS. These exposure mechanisms serve as determinants of user behaviors, which, in turn, are subsequently utilized as training data for RS (Mansoury et al., 2020; Wang et al., 2023b). Disregarding the aforementioned biases and unquestioningly adhering to the original training strategy would not only introduce additional inherent biases but also lead to the emergence of a ”rich get richer” Matthew effect over time (Gao et al., 2023a).
To address this issue, some studies attempt to explore debiased learning by adversarial training (Krishnan et al., 2018; Anelli et al., 2021), causal modeling (Zheng et al., 2021; Wu et al., 2019d) and negative sampling (Ding et al., 2019; Chen et al., 2023b). It is able to improve the fairness of RS by rectifying the impact of biases, whether they stem from data imbalances, user historical behaviors, or various other factors (Chen et al., 2023a). Existing negative sampling algorithms primarily concentrate on mitigating exposure bias, popularity bias, and unfairness, which is achieved by over-sampling popular items (Gantner et al., 2012; Cheng et al., 2021; He et al., 2016) or enhancing the sampler with side information to facilitate model debiasing (Chen et al., 2023b; Steck, 2018; Zhao et al., 2019). Nevertheless, they have not proposed solutions for the broader issues of selection bias, conformity bias, and other biases in RS. The adaptive negative sampling strategies ensure that the learning process accounts for data imbalances and mitigates any potential biases in RS. Investigating how various negative sampling strategies affect the inherent biases in recommenders would be a genuinely interesting and valuable endeavor.
5.5. Incorporating Large Language Model into Recommendation
The rapid developments of Large Language Model (LLM) in Natural Language Processing have recently gained significant attention and revolutionized the learning paradigm of diverse research fields (Huang and Chen-Chuan Chang, 2023; Zhao et al., 2023; Fang et al., 2024; Jiang et al., 2025). Some studies have indicated that the model size of LLM continues to expand with the immense scale of the corpus size, concomitantly enhancing their capabilities in logical reasoning and generalization (Hadi et al., 2023). Benefits from the potential to bridge the gap between the down-streaming task and the open-world knowledge, LLM have been incorporated into recommendation to serve as the knowledge extractor (Wu et al., 2021b; Hou et al., 2023, 2022) and the recommender (Dai et al., 2023; Gao et al., 2023b; Hou et al., 2024; Kang et al., 2023; Liu et al., 2023b; Ma et al., 2025). Most of the existing LLM-based recommendation algorithms typically employ random sampling (Bao et al., 2025; hao et al., 2023; Mysore et al., 2023; Geng et al., 2022; Du et al., 2024; Wu et al., 2024d) or leverage users’ explicit scores (Lin et al., 2024; Xi et al., 2024; Dai et al., 2023; Kang et al., 2023; Li et al., 2023; Lin et al., 2024) in real-world RS (Ratings within the range of are regarded as positive feedback, while defining the rest as negative feedback) for training. The negative samples selected in M6Rec (Cui et al., 2022) are exclusively from real-world systems, that is, Taobao and Alipay. To prevent the excessive training time and the overwhelming influx of negative information that could impact the model’s training process, M6Rec further down-samples the negative samples by a factor of ten. PALR (Yang et al., 2023a) incorporates the negative samples which are similar to the positive samples (e.g., movies belonging to the same genres and have been co-watched by many users) into model optimization.
From the comprehensive review of existing LLM-based recommendation studies, we notice that LLMRank (Hou et al., 2024) first investigates the ranking performance of LLM on hard negative candidates, which are retrieved by different classical recommendation algorithms. Moreover, InstructRec (Zhang et al., 2023a) has also delved into LLM’s effectiveness and robustness in tasks involving more challenging hard negative samples and larger candidate sets. Nevertheless, these LLM-based recommendation algorithms have not taken into account negative sampling. Furthermore, the prevalent paradigm in these algorithms lies in employing natural language for depicting user interests, context, task formats, and historical behaviors. However, the omission of users’ negative preferences in this paradigm introduces the potential bias during the training and inference of LLM, thereby impacting its downstream recommendation performance. A lot of research has demonstrated that the adoption of well-reasoned negative sampling techniques in RS can further bolster the recommenders’ performance (Shi et al., 2023; Zhou et al., 2021; Lian et al., 2020a; Wu et al., 2021a). Negative Sampling would serve as a bridge between LLM and recommendation algorithms, facilitating the smooth incorporation of LLM into recommendation tasks. This contributes to being a powerful tool for assisting future researchers in constructing a LLM-based general recommendation framework.
5.6. Understanding Negative Sampling with Theoretical Tools
To alleviate the inherent sparsity and imbalance in RS, early studies attempt to incorporate suitable negative samples into training from the corpus (auxiliary information or similarity ranking) to provide unbiased negative signals. These works simply attribute their superior performance to the improved convergence of the informative negative samples, without delving further into the profound reasons behind it. Recently, some studies have sought to understand the benefits brought to recommender by negative samples (Petrov and Macdonald, 2023) and hard negative samples (Shi et al., 2023) through the several theoretical tools. Specifically, gSASRec (Petrov and Macdonald, 2023) regards the overconfidence phenomenon of negative sampling as it increases the proportion of positive samples in the training data distribution to force recommenders to overestimate the probabilities of future user-item interactions. It not only results in the recommender’s inability to capture subtle variations within high-scored items but also hinders its effectiveness and training convergence. To address this, gSASRec theoretically defines overconfidence through a probabilistic interpretation of sequential recommendation and proposes gBCE loss to mitigate this issue via negative sampling. Inspired by the existing recognition that uniformly sampled negatives may contribute little to the gradients and model convergence (Shi et al., 2023; Rendle and Freudenthaler, 2014; Zhang et al., 2013; Ma et al., 2023), DNS∗ (Shi et al., 2023) proves that fast convergence may not be the only advantage of HNS by comparing existing hard negative sampling methods (See Sec.3.2.1) with the Non-sampling strategies (See Sec.3.1.4). Then DNS∗ establish the theoretical foundations and explain the performance gain of HNS by conducting theoretical analyses and simulation studies on One-way Partial AUC. In recent times, these works have demonstrated the superiority of utilizing theoretical tools to understand and guide negative sampling. This proves instrumental in facilitating a deeper exploration of the essence of negative samples in RS and uncovering potential opportunities for performance improvement. It stands out as a promising avenue for future research.
5.7. Exploring Sampled Metrics in Recommendation Evaluation
Personalized RS aims to leverage users’ historical behaviors to capture their interests and provide appropriate items. It is crucial to evaluate the efficacy of these recommenders during their optimization process. In large-scale RS applications, the candidate corpus often comprises millions of items, while it encompasses tens of thousands of items for academic research. Retrieving suitable samples from such an immense corpus poses a considerable challenge. To simplify the computational complexity of offline evaluations, many researchers typically employ uniform random sampling and popularity sampling in evaluating these recommendation algorithms (Ma et al., 2023, 2024b). Some recent LLM-based works even verify their effectiveness and robustness in a more challenging setting, that is, evaluating them on hard negative candidates, which are retrieved by different classical recommenders. However, some recent works have proved that these sampled metrics are inconsistent with the full item coverage in algorithm evaluation (Krichene and Rendle, 2020; Dallmann et al., 2021; Pellegrini et al., 2022; Bauer et al., 2023).
(Bauer et al., 2023) is a systematic literature review on the processes surrounding the evaluation of recommendation algorithms, which demonstrates that the predominant evaluation in recommendation algorithms is offline evaluation and that online evaluations are primarily used in combination with offline evaluation. Existing related research has explored the evaluation setting in RS from various perspectives. Precisely, (Krichene and Rendle, 2020) first explores the implications of sampling during recommender evaluation and discovers the bias in evaluation between random sampling and full item coverage. To alleviate this issue and improve the quality of the sampled metrics, a point-wise correction to the sampled metric is proposed by minimizing different criteria (e.g., bias and mean squared error). (Dallmann et al., 2021) trains four recommenders on five widely known datasets with three evaluation strategies, including full item coverage, popularity-based sampling, and random sampling. It also verifies that the evaluation results obtained from popularity-based sampling do not equal the full item coverage, which is similar to random sampling. (Pellegrini et al., 2022) investigates the correlation between optimizing popularity-sampled metrics and estimating point-wise mutual information (PMI) and proposes two PMI fitting techniques to improve popularity-sampled metrics for recommenders. How to precisely evaluate the recommendation algorithms with less computational complexity need to be given more attention for the future research.
Acknowledgements.
This work is supported in part by the TaiShan Scholars Program (Grant no. tsqn202211289), the Shandong Province Excellent Young Scientists Fund Program (Overseas) (Grant no. 2022HWYQ-048), the Oversea Innovation Team Project of the ”20 Regulations for New Universities” funding program of Jinan (Grant no. 2021GXRC073) and the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).References
- (1)
- Ali et al. (2020) Zafar Ali, Guilin Qi, Khan Muhammad, Bahadar Ali, and Waheed Ahmed Abro. 2020. Paper recommendation based on heterogeneous network embedding. Knowledge-Based Systems 210 (2020), 106438.
- Anelli et al. (2021) Vito Walter Anelli, Tommaso Di Noia, and Felice Antonio Merra. 2021. The idiosyncratic effects of adversarial training on bias in personalized recommendation learning. In Proceedings of the 15th ACM Conference on Recommender Systems. 730–735.
- Bao et al. (2025) Keqin Bao, Jizhi Zhang, Wenjie Wang, Yang Zhang, Zhengyi Yang, Yanchen Luo, Chong Chen, Fuli Feng, and Qi Tian. 2025. A Bi-Step Grounding Paradigm for Large Language Models in Recommendation Systems. ACM Trans. Recomm. Syst. 3, 4 (2025), 27 pages. https://doi.org/10.1145/3716393
- Bauer et al. (2023) Christine Bauer, Eva Zangerle, and Alan Said. 2023. Exploring the Landscape of Recommender Systems Evaluation: Practices and Perspectives. ACM Transactions on Recommender Systems (2023).
- Bengio et al. (2009) Yoshua Bengio, Jérôme Louradour, Ronan Collobert, and Jason Weston. 2009. Curriculum learning. In Proceedings of the 26th annual international conference on machine learning. 41–48.
- Bi et al. (2020) Ye Bi, Liqiang Song, Mengqiu Yao, Zhenyu Wu, Jianming Wang, and Jing Xiao. 2020. DCDIR: A deep cross-domain recommendation system for cold start users in insurance domain. In Proceedings of the 43rd international ACM SIGIR conference on research and development in information retrieval. 1661–1664.
- Brin and Page (1998) Sergey Brin and Lawrence Page. 1998. The anatomy of a large-scale hypertextual web search engine. Computer networks and ISDN systems (1998).
- Cai et al. (2023) Borui Cai, Yong Xiang, Longxiang Gao, He Zhang, Yunfeng Li, and Jianxin Li. 2023. Temporal knowledge graph completion: a survey. In Proceedings of the Thirty-Second International Joint Conference on Artificial Intelligence (IJCAI ’23). https://doi.org/10.24963/ijcai.2023/734
- Cao et al. (2022) Jiangxia Cao, Xin Cong, Jiawei Sheng, Tingwen Liu, and Bin Wang. 2022. Contrastive Cross-Domain Sequential Recommendation. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management. 138–147.
- Chen et al. (2023c) Chong Chen, Weizhi Ma, Min Zhang, Chenyang Wang, Yiqun Liu, and Shaoping Ma. 2023c. Revisiting negative sampling vs. non-sampling in implicit recommendation. ACM Transactions on Information Systems 41, 1 (2023), 1–25.
- Chen et al. (2020b) Chong Chen, Min Zhang, Weizhi Ma, Yiqun Liu, and Shaoping Ma. 2020b. Efficient non-sampling factorization machines for optimal context-aware recommendation. In Proceedings of the web conference 2020. 2400–2410.
- Chen et al. (2019d) Chong Chen, Min Zhang, Chenyang Wang, Weizhi Ma, Minming Li, Yiqun Liu, and Shaoping Ma. 2019d. An Efficient Adaptive Transfer Neural Network for Social-Aware Recommendation. In Proceedings of the 42nd International ACM SIGIR Conference on Research and Development in Information Retrieval.
- Chen et al. (2020c) Chong Chen, Min Zhang, Yongfeng Zhang, Yiqun Liu, and Shaoping Ma. 2020c. Efficient neural matrix factorization without sampling for recommendation. ACM Transactions on Information Systems (TOIS) 38, 2 (2020), 1–28.
- Chen et al. (2020d) Chong Chen, Min Zhang, Yongfeng Zhang, Weizhi Ma, Yiqun Liu, and Shaoping Ma. 2020d. Efficient heterogeneous collaborative filtering without negative sampling for recommendation. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 34. 19–26.
- Chen et al. (2020a) Fenxiao Chen, Yun-Cheng Wang, Bin Wang, and C-C Jay Kuo. 2020a. Graph representation learning: a survey. APSIPA Transactions on Signal and Information Processing 9 (2020), e15.
- Chen et al. (2021b) Hong Chen, Yudong Chen, Xin Wang, Ruobing Xie, Rui Wang, Feng Xia, and Wenwu Zhu. 2021b. Curriculum disentangled recommendation with noisy multi-feedback. Advances in Neural Information Processing Systems 34 (2021), 26924–26936.
- Chen et al. (2023d) Hong Chen, Xin Wang, Ruobing Xie, Yuwei Zhou, and Wenwu Zhu. 2023d. Cross-domain Recommendation with Behavioral Importance Perception. In Proceedings of the World Wide Web Conference (WWW).
- Chen et al. (2021c) Jiawei Chen, Hande Dong, Yang Qiu, Xiangnan He, Xin Xin, Liang Chen, Guli Lin, and Keping Yang. 2021c. AutoDebias: Learning to debias for recommendation. In Proceedings of the 44th International ACM SIGIR Conference on Research and Development in Information Retrieval. 21–30.
- Chen et al. (2023a) Jiawei Chen, Hande Dong, Xiang Wang, Fuli Feng, Meng Wang, and Xiangnan He. 2023a. Bias and debias in recommender system: A survey and future directions. ACM Transactions on Information Systems 41, 3 (2023), 1–39.
- Chen et al. (2022a) Jingfan Chen, Wenqi Fan, Guanghui Zhu, Xiangyu Zhao, Chunfeng Yuan, Qing Li, and Yihua Huang. 2022a. Knowledge-enhanced Black-box Attacks for Recommendations. In Proceedings of the 28th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. 108–117.
- Chen et al. (2021d) Junyang Chen, Zhiguo Gong, Wei Wang, and Weiwen Liu. 2021d. HNS: Hierarchical negative sampling for network representation learning. Information Sciences 542 (2021), 343–356.
- Chen et al. (2019b) Jiawei Chen, Can Wang, Sheng Zhou, Qihao Shi, Yan Feng, and Chun Chen. 2019b. SamWalker: Social Recommendation with Informative Sampling Strategy. In Proceedings of the World Wide Web Conference (WWW).
- Chen et al. (2023e) Jiawei Chen, Junkang Wu, Jiancan Wu, Xuezhi Cao, Sheng Zhou, and Xiangnan He. 2023e. Adap-: Adaptively Modulating Embedding Magnitude for Recommendation. In Proceedings of the ACM Web Conference 2023 (Austin, TX, USA) (WWW ’23). 1085–1096. https://doi.org/10.1145/3543507.3583363
- Chen et al. (2019c) Li Chen, Yonghua Yang, Ningxia Wang, Keping Yang, and Quan Yuan. 2019c. How serendipity improves user satisfaction with recommendations? a large-scale user evaluation. In The world wide web conference. 240–250.
- Chen et al. (2019a) Minmin Chen, Alex Beutel, Paul Covington, Sagar Jain, Francois Belletti, and Ed H Chi. 2019a. Top-k off-policy correction for a REINFORCE recommender system. In Proceedings of the Twelfth ACM International Conference on Web Search and Data Mining. 456–464.
- Chen et al. (2021a) Minmin Chen, Bo Chang, Can Xu, and Ed H Chi. 2021a. User response models to improve a reinforce recommender system. In Proceedings of the 14th ACM International Conference on Web Search and Data Mining. 121–129.
- Chen et al. (2016) Tao Chen, Xiangnan He, and Min-Yen Kan. 2016. Context-aware image tweet modelling and recommendation. In Proceedings of the 24th ACM international conference on Multimedia. 1018–1027.
- Chen et al. (2017) Ting Chen, Yizhou Sun, Yue Shi, and Liangjie Hong. 2017. On sampling strategies for neural network-based collaborative filtering. In Proceedings of the 23rd ACM SIGKDD International Conference on Knowledge Discovery and Data Mining. 767–776.
- Chen et al. (2022c) Wei Chen, Yu Liu, Weiping Wang, Erwin M Bakker, Theodoros Georgiou, Paul Fieguth, Li Liu, and Michael S Lew. 2022c. Deep learning for instance retrieval: A survey. IEEE Transactions on Pattern Analysis and Machine Intelligence (2022).
- Chen et al. (2023b) Xiao Chen, Wenqi Fan, Jingfan Chen, Haochen Liu, Zitao Liu, Zhaoxiang Zhang, and Qing Li. 2023b. Fairly Adaptive Negative Sampling for Recommendations. In Proceedings of the ACM Web Conference 2023 (Austin, TX, USA) (WWW ’23). 3723–3733. https://doi.org/10.1145/3543507.3583355
- Chen et al. (2022b) Yongjun Chen, Jia Li, and Caiming Xiong. 2022b. ELECRec: Training Sequential Recommenders as Discriminators. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval. 2550–2554.
- Chen et al. (2024) Yuxin Chen, Junfei Tan, An Zhang, Zhengyi Yang, Leheng Sheng, Enzhi Zhang, Xiang Wang, and Tat-Seng Chua. 2024. On Softmax Direct Preference Optimization for Recommendation. In Advances in Neural Information Processing Systems, A. Globerson, L. Mackey, D. Belgrave, A. Fan, U. Paquet, J. Tomczak, and C. Zhang (Eds.), Vol. 37. Curran Associates, Inc., 27463–27489. https://proceedings.neurips.cc/paper_files/paper/2024/file/30732ddb12d9faf7180f5d0e8b5b5da7-Paper-Conference.pdf
- Chen et al. (2021e) Yudong Chen, Xin Wang, Miao Fan, Jizhou Huang, Shengwen Yang, and Wenwu Zhu. 2021e. Curriculum meta-learning for next POI recommendation. In Proceedings of the 27th ACM SIGKDD Conference on Knowledge Discovery & Data Mining.
- Cheng et al. (2021) Mingyue Cheng, Fajie Yuan, Qi Liu, Shenyang Ge, Zhi Li, Runlong Yu, Defu Lian, Senchao Yuan, and Enhong Chen. 2021. Learning recommender systems with implicit feedback via soft target enhancement. In Proceedings of the 44th International ACM SIGIR Conference on Research and Development in Information Retrieval. 575–584.
- Covington et al. (2016) Paul Covington, Jay Adams, and Emre Sargin. 2016. Deep neural networks for youtube recommendations. In Proceedings of the 10th ACM conference on recommender systems. 191–198.
- Cui et al. (2022) Zeyu Cui, Jianxin Ma, Chang Zhou, Jingren Zhou, and Hongxia Yang. 2022. M6-Rec: Generative Pretrained Language Models are Open-Ended Recommender Systems. Association for Computing Machinery. http://arxiv.org/abs/2205.08084
- Dai et al. (2024a) Sunhao Dai, Changle Qu, Sirui Chen, Xiao Zhang, and Jun Xu. 2024a. ReCODE: Modeling Repeat Consumption with Neural ODE. In Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval. 2599–2603.
- Dai et al. (2023) Sunhao Dai, Ninglu Shao, Haiyuan Zhao, Weijie Yu, Zihua Si, Chen Xu, Zhongxiang Sun, Xiao Zhang, and Jun Xu. 2023. Uncovering ChatGPTś Capabilities in Recommender Systems. (2023). https://doi.org/10.1145/3604915.3610646
- Dai et al. (2024b) Sunhao Dai, Ninglu Shao, Jieming Zhu, Xiao Zhang, Zhenhua Dong, Jun Xu, Quanyu Dai, and Ji-Rong Wen. 2024b. Modeling User Attention in Music Recommendation. In 2024 IEEE 40th International Conference on Data Engineering (ICDE). IEEE, 761–774.
- Dallmann et al. (2021) Alexander Dallmann, Daniel Zoller, and Andreas Hotho. 2021. A case study on sampling strategies for evaluating neural sequential item recommendation models. In Proceedings of the 15th ACM Conference on Recommender Systems. 505–514.
- Ding et al. (2019) Jingtao Ding, Yuhan Quan, Xiangnan He, Yong Li, and Depeng Jin. 2019. Reinforced Negative Sampling for Recommendation with Exposure Data.. In IJCAI. Macao, 2230–2236.
- Ding et al. (2020) Jingtao Ding, Yuhan Quan, Quanming Yao, Yong Li, and Depeng Jin. 2020. Simplify and robustify negative sampling for implicit collaborative filtering. Advances in Neural Information Processing Systems 33 (2020), 1094–1105.
- Ding et al. (2023) Rui Ding, Ruobing Xie, Xiaobo Hao, Xiaochun Yang, Kaikai Ge, Xu Zhang, Jie Zhou, and Leyu Lin. 2023. Interpretable User Retention Modeling in Recommendation. In Proceedings of the 17th ACM Conference on Recommender Systems. 702–708.
- Dong et al. (2024) Xue Dong, Xuemeng Song, Minghui Tian, and Linmei Hu. 2024. Prompt-based and weak-modality enhanced multimodal recommendation. Information Fusion 101 (2024).
- Du et al. (2022) Hanwen Du, Hui Shi, Pengpeng Zhao, Deqing Wang, Victor S Sheng, Yanchi Liu, Guanfeng Liu, and Lei Zhao. 2022. Contrastive Learning with Bidirectional Transformers for Sequential Recommendation. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management. 396–405.
- Du et al. (2024) Yingpeng Du, Di Luo, Rui Yan, Xiaopei Wang, Hongzhi Liu, Hengshu Zhu, Yang Song, and Jie Zhang. 2024. Enhancing job recommendation through llm-based generative adversarial networks. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38. 8363–8371.
- Fan (2024) Wenqi Fan. 2024. Recommender systems in the era of large language models (llms). IEEE Transactions on Knowledge and Data Engineering (2024), 1–20.
- Fang et al. (2024) Junfeng Fang, Houcheng Jiang, Kun Wang, Yunshan Ma, Xiang Wang, Xiangnan He, and Tat-seng Chua. 2024. Alphaedit: Null-space constrained knowledge editing for language models. arXiv preprint arXiv:2410.02355 (2024).
- Gantner et al. (2012) Zeno Gantner, Lucas Drumond, Christoph Freudenthaler, and Lars Schmidt-Thieme. 2012. Personalized ranking for non-uniformly sampled items. In Proceedings of KDD Cup 2011. PMLR, 231–247.
- Gao et al. (2023a) Chongming Gao, Kexin Huang, Jiawei Chen, Yuan Zhang, Biao Li, Peng Jiang, Shiqi Wang, Zhong Zhang, and Xiangnan He. 2023a. Alleviating matthew effect of offline reinforcement learning in interactive recommendation. In Proceedings of the 46th international ACM SIGIR conference on research and development in information retrieval. 238–248.
- Gao et al. (2023c) Chen Gao, Yu Zheng, Nian Li, Yinfeng Li, Yingrong Qin, Jinghua Piao, Yuhan Quan, Jianxin Chang, Depeng Jin, Xiangnan He, et al. 2023c. A survey of graph neural networks for recommender systems: Challenges, methods, and directions. ACM Transactions on Recommender Systems (2023).
- Gao et al. (2024) Chen Gao, Yu Zheng, Wenjie Wang, Fuli Feng, Xiangnan He, and Yong Li. 2024. Causal inference in recommender systems: A survey and future directions. ACM Transactions on Information Systems 42, 4 (2024), 1–32.
- Gao et al. (2019) Rong Gao, Haifeng Xia, Jing Li, Donghua Liu, Shuai Chen, and Gang Chun. 2019. DRCGR: Deep reinforcement learning framework incorporating CNN and GAN-based for interactive recommendation. In 2019 IEEE International Conference on Data Mining (ICDM). IEEE, 1048–1053.
- Gao et al. (2022) Yunjun Gao, Yuntao Du, Yujia Hu, Lu Chen, Xinjun Zhu, Ziquan Fang, and Baihua Zheng. 2022. Self-guided learning to denoise for robust recommendation. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval. 1412–1422.
- Gao et al. (2023b) Yunfan Gao, Tao Sheng, Youlin Xiang, Yun Xiong, Haofen Wang, and Jiawei Zhang. 2023b. Chat-REC: Towards Interactive and Explainable LLMs-Augmented Recommender System. (2023). http://arxiv.org/abs/2303.14524
- Geng et al. (2022) Shijie Geng, Shuchang Liu, Zuohui Fu, Yingqiang Ge, and Yongfeng Zhang. 2022. Recommendation as Language Processing (RLP): A Unified Pretrain, Personalized Prompt & Predict Paradigm (P5). (2022). https://doi.org/10.1145/3523227.3546767
- Giobergia (2022) Flavio Giobergia. 2022. Triplet Losses-based Matrix Factorization for Robust Recommendations. CoRR abs/2210.12098 (2022). https://doi.org/10.48550/ARXIV.2210.12098 arXiv:2210.12098
- Gong et al. (2022) Kaiqi Gong, Xiao Song, Senzhang Wang, Songsong Liu, and Yong Li. 2022. ITSM-GCN: Informative Training Sample Mining for Graph Convolutional Network-based Collaborative Filtering. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management. 614–623.
- Goodfellow et al. (2014) Ian J. Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron Courville, and Yoshua Bengio. 2014. Generative Adversarial Networks.
- Gu et al. (2022) Shuyun Gu, Xiao Wang, Chuan Shi, and Ding Xiao. 2022. Self-supervised Graph Neural Networks for Multi-behavior Recommendation. In International Joint Conference on Artificial Intelligence (IJCAI).
- Gu et al. (2021) Yulong Gu, Wentian Bao, Dan Ou, Xiang Li, Baoliang Cui, Biyu Ma, Haikuan Huang, Qingwen Liu, and Xiaoyi Zeng. 2021. Self-Supervised Learning on Users’ Spontaneous Behaviors for Multi-Scenario Ranking in E-commerce. In Proceedings of the 30th ACM International Conference on Information & Knowledge Management. 3828–3837.
- Guo et al. (2022) Jiafeng Guo, Yinqiong Cai, Yixing Fan, Fei Sun, Ruqing Zhang, and Xueqi Cheng. 2022. Semantic models for the first-stage retrieval: A comprehensive review. ACM Transactions on Information Systems (TOIS) 40, 4 (2022), 1–42.
- Guo et al. (2020a) Lei Guo, Hongzhi Yin, Qinyong Wang, Bin Cui, Zi Huang, and Lizhen Cui. 2020a. Group recommendation with latent voting mechanism. In 2020 IEEE 36th International Conference on Data Engineering (ICDE). IEEE, 121–132.
- Guo et al. (2020b) Qingyu Guo, Fuzhen Zhuang, Chuan Qin, Hengshu Zhu, Xing Xie, Hui Xiong, and Qing He. 2020b. A survey on knowledge graph-based recommender systems. IEEE Transactions on Knowledge and Data Engineering 34, 8 (2020), 3549–3568.
- Ha et al. (2024) Mingming Ha, Xuewen Tao, Wenfang Lin, Qiongxu Ma, Wujiang Xu, and Linxun Chen. 2024. Fine-Grained Dynamic Framework for Bias-Variance Joint Optimization on Data Missing Not at Random. Advances in Neural Information Processing Systems 37 (2024), 104010–104034.
- Hadi et al. (2023) Muhammad Usman Hadi, Rizwan Qureshi, Abbas Shah, Muhammad Irfan, Anas Zafar, Muhammad Bilal Shaikh, Naveed Akhtar, Jia Wu, Seyedali Mirjalili, et al. 2023. Large language models: a comprehensive survey of its applications, challenges, limitations, and future prospects. (2023).
- Han et al. (2021) Junlin Han, Mehrdad Shoeiby, Lars Petersson, and Mohammad Ali Armin. 2021. Dual contrastive learning for unsupervised image-to-image translation. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 746–755.
- hao et al. (2023) Chen hao, Xie Runfeng, Cui Xiangyang, Yan Zhou, Wang Xin, Xuan Zhanwei, and Zhang Kai. 2023. LKPNR: LLM and KG for Personalized News Recommendation Framework. arXiv:2308.12028 [cs.IR]
- Hao et al. (2021b) Qianxiu Hao, Qianqian Xu, Zhiyong Yang, and Qingming Huang. 2021b. Learning unified embeddings for recommendation via meta-path semantics. In Proceedings of the 29th ACM International Conference on Multimedia. 3909–3917.
- Hao et al. (2021a) Xiaobo Hao, Yudan Liu, Ruobing Xie, Kaikai Ge, Linyao Tang, Xu Zhang, and Leyu Lin. 2021a. Adversarial feature translation for multi-domain recommendation. In Proceedings of the 27th ACM SIGKDD Conference on Knowledge Discovery & Data Mining. 2964–2973.
- He et al. (2023a) Li He, Xianzhi Wang, Dingxian Wang, Haoyuan Zou, Hongzhi Yin, and Guandong Xu. 2023a. Simplifying Graph-based Collaborative Filtering for Recommendation. In Proceedings of the Sixteenth ACM International Conference on Web Search and Data Mining. 60–68.
- He et al. (2017) Xiangnan He, Lizi Liao, Hanwang Zhang, Liqiang Nie, Xia Hu, and Tat-Seng Chua. 2017. Neural collaborative filtering. In Proceedings of the 26th international conference on world wide web. 173–182.
- He et al. (2016) Xiangnan He, Hanwang Zhang, Min-Yen Kan, and Tat-Seng Chua. 2016. Fast matrix factorization for online recommendation with implicit feedback. In Proceedings of the 39th International ACM SIGIR conference on Research and Development in Information Retrieval. 549–558.
- He et al. (2023b) Xiangnan He, Yang Zhang, Fuli Feng, Chonggang Song, Lingling Yi, Guohui Ling, and Yongdong Zhang. 2023b. Addressing confounding feature issue for causal recommendation. ACM Transactions on Information Systems 41, 3 (2023), 1–23.
- He et al. (2022) Yue He, Zimu Wang, Peng Cui, Hao Zou, Yafeng Zhang, Qiang Cui, and Yong Jiang. 2022. Causpref: Causal preference learning for out-of-distribution recommendation. In Proceedings of the ACM Web Conference 2022. 410–421.
- Hidasi and Karatzoglou (2018) Balázs Hidasi and Alexandros Karatzoglou. 2018. Recurrent neural networks with top-k gains for session-based recommendations. In Proceedings of the 27th ACM international conference on information and knowledge management. 843–852.
- Hidasi et al. (2015) Balázs Hidasi, Alexandros Karatzoglou, Linas Baltrunas, and Domonkos Tikk. 2015. Session-based recommendations with recurrent neural networks. arXiv preprint arXiv:1511.06939 (2015).
- Hou et al. (2023) Yupeng Hou, Zhankui He, Julian McAuley, and Wayne Xin Zhao. 2023. Learning vector-quantized item representation for transferable sequential recommenders. In Proceedings of the ACM Web Conference 2023.
- Hou et al. (2022) Yupeng Hou, Shanlei Mu, Wayne Xin Zhao, Yaliang Li, Bolin Ding, and Ji-Rong Wen. 2022. Towards universal sequence representation learning for recommender systems. In Proceedings of the 28th ACM SIGKDD Conference on Knowledge Discovery and Data Mining.
- Hou et al. (2024) Yupeng Hou, Junjie Zhang, Zihan Lin, Hongyu Lu, Ruobing Xie, Julian McAuley, and Wayne Xin Zhao. 2024. Large language models are zero-shot rankers for recommender systems. In European Conference on Information Retrieval. Springer, 364–381.
- Hu et al. (2018) Binbin Hu, Chuan Shi, Wayne Xin Zhao, and Philip S Yu. 2018. Leveraging meta-path based context for top-n recommendation with a neural co-attention model. In Proceedings of the 24th ACM SIGKDD international conference on knowledge discovery & data mining. 1531–1540.
- Huang and Chen-Chuan Chang (2023) Jie Huang and Kevin Chen-Chuan Chang. 2023. Towards Reasoning in Large Language Models: A Survey. In Findings of the Association for Computational Linguistics, ACL 2023 (Proceedings of the Annual Meeting of the Association for Computational Linguistics). Association for Computational Linguistics (ACL), 1049–1065. https://doi.org/10.18653/v1/2023.findings-acl.67
- Huang et al. (2023) Junjie Huang, Ruobing Xie, Qi Cao, Huawei Shen, Shaoliang Zhang, Feng Xia, and Xueqi Cheng. 2023. Negative can be positive: Signed graph neural networks for recommendation. Information Processing & Management 60, 4 (2023), 103403.
- Huang et al. (2020) Jui-Ting Huang, Ashish Sharma, Shuying Sun, Li Xia, David Zhang, Philip Pronin, Janani Padmanabhan, Giuseppe Ottaviano, and Linjun Yang. 2020. Embedding-based retrieval in facebook search. In Proceedings of the 26th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining. 2553–2561.
- Huang et al. (2021) Tinglin Huang, Yuxiao Dong, Ming Ding, Zhen Yang, Wenzheng Feng, Xinyu Wang, and Jie Tang. 2021. Mixgcf: An improved training method for graph neural network-based recommender systems. In Proceedings of the 27th ACM SIGKDD Conference on Knowledge Discovery & Data Mining.
- Jiang et al. (2025) Houcheng Jiang, Junfeng Fang, Ningyu Zhang, Guojun Ma, Mingyang Wan, Xiang Wang, Xiangnan He, and Tat-seng Chua. 2025. AnyEdit: Edit Any Knowledge Encoded in Language Models. arXiv preprint arXiv:2502.05628 (2025).
- Kamigaito and Hayashi (2022) Hidetaka Kamigaito and Katsuhiko Hayashi. 2022. Comprehensive analysis of negative sampling in knowledge graph representation learning. In International Conference on Machine Learning. PMLR, 10661–10675.
- Kang et al. (2020) SeongKu Kang, Junyoung Hwang, Wonbin Kweon, and Hwanjo Yu. 2020. DE-RRD: A knowledge distillation framework for recommender system. In Proceedings of the 29th ACM International Conference on Information & Knowledge Management. 605–614.
- Kang et al. (2023) Wang-Cheng Kang, Jianmo Ni, Nikhil Mehta, Maheswaran Sathiamoorthy, Lichan Hong, Ed Chi, and Derek Zhiyuan Cheng. 2023. Do LLMs Understand User Preferences? Evaluating LLMs On User Rating Prediction. (2023). http://arxiv.org/abs/2305.06474
- Kaya and Bilge (2019) Mahmut Kaya and Hasan Şakir Bilge. 2019. Deep metric learning: A survey. Symmetry 11, 9 (2019), 1066.
- Khoshraftar and An (2022) Shima Khoshraftar and Aijun An. 2022. A survey on graph representation learning methods. arXiv preprint arXiv:2204.01855 (2022).
- Kim et al. (2021a) Minseok Kim, Hwanjun Song, Doyoung Kim, Kijung Shin, and Jae-Gil Lee. 2021a. Premere: Meta-reweighting via self-ensembling for point-of-interest recommendation. In Proceedings of the AAAI Conference on artificial intelligence, Vol. 35. 4164–4171.
- Kim et al. (2021b) Taeuk Kim, Kang Min Yoo, and Sang goo Lee. 2021b. Self-Guided Contrastive Learning for BERT Sentence Representations. ArXiv abs/2106.07345 (2021). https://api.semanticscholar.org/CorpusID:235422673
- Krichene and Rendle (2020) Walid Krichene and Steffen Rendle. 2020. On sampled metrics for item recommendation. In Proceedings of the 26th ACM SIGKDD international conference on knowledge discovery & data mining. 1748–1757.
- Krishnan et al. (2020) Adit Krishnan, Mahashweta Das, Mangesh Bendre, Hao Yang, and Hari Sundaram. 2020. Transfer learning via contextual invariants for one-to-many cross-domain recommendation. In Proceedings of the 43rd International ACM SIGIR Conference on Research and Development in Information Retrieval. 1081–1090.
- Krishnan et al. (2018) Adit Krishnan, Ashish Sharma, Aravind Sankar, and Hari Sundaram. 2018. An adversarial approach to improve long-tail performance in neural collaborative filtering. In Proceedings of the 27th ACM International Conference on information and knowledge management. 1491–1494.
- Lai et al. (2023) Riwei Lai, Li Chen, Yuhan Zhao, Rui Chen, and Qilong Han. 2023. Disentangled Negative Sampling for Collaborative Filtering. In Proceedings of the Sixteenth ACM International Conference on Web Search and Data Mining. 96–104.
- Lai et al. (2024) Riwei Lai, Rui Chen, Qilong Han, Chi Zhang, and Li Chen. 2024. Adaptive hardness negative sampling for collaborative filtering. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 38. 8645–8652.
- Lee et al. (2019) Jae-woong Lee, Minjin Choi, Jongwuk Lee, and Hyunjung Shim. 2019. Collaborative distillation for top-N recommendation. In 2019 IEEE International Conference on Data Mining (ICDM). IEEE, 369–378.
- Lei et al. (2020) Wenqiang Lei, Xiangnan He, Yisong Miao, Qingyun Wu, Richang Hong, Min-Yen Kan, and Tat-Seng Chua. 2020. Estimation-action-reflection: Towards deep interaction between conversational and recommender systems. In Proceedings of the 13th International Conference on Web Search and Data Mining. 304–312.
- Li et al. (2024) Haoxuan Li, Chunyuan Zheng, Wenjie Wang, Hao Wang, Fuli Feng, and Xiao-Hua Zhou. 2024. Debiased recommendation with noisy feedback. In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. 1576–1586.
- Li et al. (2022) Nian Li, Chen Gao, Jinghua Piao, Xin Huang, Aizhen Yue, Liang Zhou, Qingmin Liao, and Yong Li. 2022. An Exploratory Study of Information Cocoon on Short-form Video Platform. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management. 4178–4182.
- Li et al. (2023) Xiangyang Li, Bo Chen, Lu Hou, and Ruiming Tang. 2023. Ctrl: Connect collaborative and language model for ctr prediction. ACM Transactions on Recommender Systems (2023).
- Li et al. (2020) Yang Li, Yadan Luo, and Zi Huang. 2020. Fashion recommendation with multi-relational representation learning. In Advances in Knowledge Discovery and Data Mining: 24th Pacific-Asia Conference, PAKDD 2020, Singapore, May 11–14, 2020, Proceedings, Part I 24. Springer, 3–15.
- Li et al. (2021) Zelong Li, Jianchao Ji, Zuohui Fu, Yingqiang Ge, Shuyuan Xu, Chong Chen, and Yongfeng Zhang. 2021. Efficient non-sampling knowledge graph embedding. In Proceedings of the Web Conference 2021. 1727–1736.
- Li et al. (2018) Zhi Li, Hongke Zhao, Qi Liu, Zhenya Huang, Tao Mei, and Enhong Chen. 2018. Learning from history and present: Next-item recommendation via discriminatively exploiting user behaviors. In Proceedings of the 24th ACM SIGKDD international conference on knowledge discovery & data mining.
- Lian et al. (2020a) Defu Lian, Qi Liu, and Enhong Chen. 2020a. Personalized ranking with importance sampling. In Proceedings of The Web Conference 2020. 1093–1103.
- Lian et al. (2020b) Defu Lian, Yongji Wu, Yong Ge, Xing Xie, and Enhong Chen. 2020b. Geography-aware sequential location recommendation. In Proceedings of the 26th ACM SIGKDD international conference on knowledge discovery & data mining. 2009–2019.
- Lian et al. (2014) Defu Lian, Cong Zhao, Xing Xie, Guangzhong Sun, Enhong Chen, and Yong Rui. 2014. GeoMF: joint geographical modeling and matrix factorization for point-of-interest recommendation. In Proceedings of the 20th ACM SIGKDD international conference on Knowledge discovery and data mining.
- Lin et al. (2024) Jianghao Lin, Rong Shan, Chenxu Zhu, Kounianhua Du, Bo Chen, Shigang Quan, Ruiming Tang, Yong Yu, and Weinan Zhang. 2024. Rella: Retrieval-enhanced large language models for lifelong sequential behavior comprehension in recommendation. Proceedings of the ACM on Web Conference.
- Liu et al. (2021b) Danyang Liu, Jianxun Lian, Zheng Liu, Xiting Wang, Guangzhong Sun, and Xing Xie. 2021b. Reinforced anchor knowledge graph generation for news recommendation reasoning. In Proceedings of the 27th ACM SIGKDD Conference on Knowledge Discovery & Data Mining. 1055–1065.
- Liu et al. (2021a) Huiting Liu, Lingling Guo, Peipei Li, Peng Zhao, and Xindong Wu. 2021a. Collaborative filtering with a deep adversarial and attention network for cross-domain recommendation. Information Sciences 565 (2021), 370–389.
- Liu et al. (2023b) Junling Liu, Chao Liu, Peilin Zhou, Renjie Lv, Kang Zhou, and Yan Zhang. 2023b. Is ChatGPT a Good Recommender? A Preliminary Study. (2023). http://arxiv.org/abs/2304.10149
- Liu et al. (2020) Meng Liu, Jianjun Li, Guohui Li, and Peng Pan. 2020. Cross domain recommendation via bi-directional transfer graph collaborative filtering networks. In Proceedings of the 29th ACM international conference on information & knowledge management. 885–894.
- Liu et al. (2023a) Xiaoyang Liu, Chong Liu, Pinzheng Wang, Rongqin Zheng, Lixin Zhang, Leyu Lin, Zhijun Chen, and Liangliang Fu. 2023a. UFNRec: Utilizing False Negative Samples for Sequential Recommendation. In Proceedings of the 2023 SIAM International Conference on Data Mining (SDM). SIAM, 46–54.
- Liu and Tang (2021) Xueyi Liu and Jie Tang. 2021. Network representation learning: A macro and micro view. AI Open 2 (2021), 43–64.
- Lou et al. (2019) Chen Lou, Quan Xie, Yang Feng, and Wonkyung Kim. 2019. Does non-hard-sell content really work? Leveraging the value of branded content marketing in brand building. Journal of Product & Brand Management 28, 7 (2019), 773–786.
- Ma et al. (2018) Chen Ma, Yingxue Zhang, Qinglong Wang, and Xue Liu. 2018. Point-of-interest recommendation: Exploiting self-attentive autoencoders with neighbor-aware influence. In Proceedings of the 27th ACM international conference on information and knowledge management. 697–706.
- Ma et al. (2025) Haokai Ma, Yunshan Ma, Ruobing Xie, Lei Meng, Jialie Shen, Xingwu Sun, Zhanhui Kang, and Tat-Seng Chua. 2025. Large Language Model Empowered Recommendation Meets All-domain Continual Pre-Training. arXiv preprint arXiv:2504.08949 (2025).
- Ma et al. (2024a) Haokai Ma, Ruobing Xie, Lei Meng, Xin Chen, Xu Zhang, Leyu Lin, and Zhanhui Kang. 2024a. Plug-in Diffusion Model for Sequential Recommendation. In Proceedings of the AAAI Conference on Artificial Intelligence.
- Ma et al. (2023) Haokai Ma, Ruobing Xie, Lei Meng, Xin Chen, Xu Zhang, Leyu Lin, and Jie Zhou. 2023. Exploring False Hard Negative Sample in Cross-Domain Recommendation. In Proceedings of the 17th ACM Conference on Recommender Systems.
- Ma et al. (2024b) Haokai Ma, Ruobing Xie, Lei Meng, Xin Chen, Xu Zhang, Leyu Lin, and Jie Zhou. 2024b. Triple sequence learning for cross-domain recommendation. ACM Transactions on Information Systems 42, 4 (2024), 1–29.
- Ma et al. (2024c) Haokai Ma, Ruobing Xie, Lei Meng, Yimeng Yang, Xingwu Sun, and Zhanhui Kang. 2024c. SeeDRec: Sememe-based Diffusion for Sequential Recommendation. In Proceedings of the Thirty-Third International Joint Conference on Artificial Intelligence, IJCAI-24. International Joint Conferences on Artificial Intelligence Organization, 2270–2278. https://doi.org/10.24963/ijcai.2024/251
- Ma et al. (2024d) Haokai Ma, Yimeng Yang, Lei Meng, Ruobing Xie, and Xiangxu Meng. 2024d. Multimodal Conditioned Diffusion Model for Recommendation. In Companion Proceedings of the ACM Web Conference 2024 (WWW ’24). 1733–1740. https://doi.org/10.1145/3589335.3651956
- Ma et al. (2022) Yunshan Ma, Yingzhi He, An Zhang, Xiang Wang, and Tat-Seng Chua. 2022. CrossCBR: cross-view contrastive learning for bundle recommendation. In Proceedings of the 28th ACM SIGKDD conference on knowledge discovery and data mining. 1233–1241.
- Mansoury et al. (2020) Masoud Mansoury, Himan Abdollahpouri, Mykola Pechenizkiy, Bamshad Mobasher, and Robin Burke. 2020. Feedback loop and bias amplification in recommender systems. In Proceedings of the 29th ACM international conference on information & knowledge management. 2145–2148.
- Meng et al. (2020) Lei Meng, Fuli Feng, Xiangnan He, Xiaoyan Gao, and Tat-Seng Chua. 2020. Heterogeneous fusion of semantic and collaborative information for visually-aware food recommendation. In Proceedings of the 28th ACM international conference on multimedia. 3460–3468.
- Mikolov et al. (2013) Tomas Mikolov, Ilya Sutskever, Kai Chen, Greg Corrado, and Jeffrey Dean. 2013. Distributed Representations of Words and Phrases and their Compositionality.
- Mo et al. (2022) Yujie Mo, Liang Peng, Jie Xu, Xiaoshuang Shi, and Xiaofeng Zhu. 2022. Simple unsupervised graph representation learning. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 36. 7797–7805.
- Mysore et al. (2023) Sheshera Mysore, Andrew Mccallum, and Hamed Zamani. 2023. Large Language Model Augmented Narrative Driven Recommendations. (2023). https://doi.org/10.1145/3604915.3608829
- Pellegrini et al. (2022) Roberto Pellegrini, Wenjie Zhao, and Iain Murray. 2022. Don’t recommend the obvious: estimate probability ratios. In Proceedings of the 16th ACM Conference on Recommender Systems. 188–197.
- Peterson (2004) Gail B Peterson. 2004. A day of great illumination: BF Skinner’s discovery of shaping. Journal of the experimental analysis of behavior 82, 3 (2004), 317–328.
- Petrov and Macdonald (2022) Aleksandr Petrov and Craig Macdonald. 2022. Effective and Efficient Training for Sequential Recommendation using Recency Sampling. In Proceedings of the 16th ACM Conference on Recommender Systems. 81–91.
- Petrov and Macdonald (2023) Aleksandr Vladimirovich Petrov and Craig Macdonald. 2023. gSASRec: Reducing Overconfidence in Sequential Recommendation Trained with Negative Sampling. In Proceedings of the 17th ACM Conference on Recommender Systems. 116–128.
- Qian et al. (2020) Jing Qian, Gangmin Li, Katie Atkinson, and Yong Yue. 2020. Negative Sampling in Knowledge Representation Learning: A Mini-Review. https://api.semanticscholar.org/CorpusID:229449355
- Qin et al. (2020) Jiarui Qin, Kan Ren, Yuchen Fang, Weinan Zhang, and Yong Yu. 2020. Sequential recommendation with dual side neighbor-based collaborative relation modeling. In Proceedings of the 13th international conference on web search and data mining. 465–473.
- Rendle and Freudenthaler (2014) Steffen Rendle and Christoph Freudenthaler. 2014. Improving pairwise learning for item recommendation from implicit feedback. In Proceedings of the 7th ACM international conference on Web search and data mining. 273–282.
- Rendle et al. (2009) Steffen Rendle, Christoph Freudenthaler, Zeno Gantner, and Lars Schmidt-Thieme. 2009. BPR: Bayesian personalized ranking from implicit feedback. In Proceedings of the Twenty-Fifth Conference on Uncertainty in Artificial Intelligence (Montreal, Quebec, Canada) (UAI ’09). AUAI Press, Arlington, Virginia, USA, 452–461.
- Saito (2020) Yuta Saito. 2020. Asymmetric tri-training for debiasing missing-not-at-random explicit feedback. In Proceedings of the 43rd International ACM SIGIR Conference on Research and Development in Information Retrieval. 309–318.
- Saito et al. (2020a) Yuta Saito, Suguru Yaginuma, Yuta Nishino, Hayato Sakata, and Kazuhide Nakata. 2020a. Unbiased recommender learning from missing-not-at-random implicit feedback. In Proceedings of the 13th International Conference on Web Search and Data Mining. 501–509.
- Saito et al. (2020b) Yuta Saito, Suguru Yaginuma, Yuta Nishino, Hayato Sakata, and Kazuhide Nakata. 2020b. Unbiased recommender learning from missing-not-at-random implicit feedback. In Proceedings of the 13th international conference on web search and data mining. 501–509.
- Shen et al. (2021) Qijie Shen, Wanjie Tao, Jing Zhang, Hong Wen, Zulong Chen, and Quan Lu. 2021. SAR-Net: A scenario-aware ranking network for personalized fair recommendation in hundreds of travel scenarios. In Proceedings of the 30th ACM International Conference on Information & Knowledge Management.
- Shi et al. (2022) Kexin Shi, Yun Zhang, Bingyi Jing, and Wenjia Wang. 2022. Soft BPR Loss for Dynamic Hard Negative Sampling in Recommender Systems. arXiv preprint arXiv:2211.13912 (2022).
- Shi et al. (2023) Wentao Shi, Jiawei Chen, Fuli Feng, Jizhi Zhang, Junkang Wu, Chongming Gao, and Xiangnan He. 2023. On the theories behind hard negative sampling for recommendation. In Proceedings of the ACM Web Conference 2023. 812–822.
- Shuai et al. (2022) Jie Shuai, Kun Zhang, Le Wu, Peijie Sun, Richang Hong, Meng Wang, and Yong Li. 2022. A review-aware graph contrastive learning framework for recommendation. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval. 1283–1293.
- Song et al. (2015) Dongjin Song, David A Meyer, and Dacheng Tao. 2015. Efficient latent link recommendation in signed networks. In Proceedings of the 21th ACM SIGKDD international conference on knowledge discovery and data mining. 1105–1114.
- Steck (2018) Harald Steck. 2018. Calibrated recommendations. In Proceedings of the 12th ACM conference on recommender systems. 154–162.
- Sun et al. (2020) Changfeng Sun, Han Liu, Meng Liu, Zhaochun Ren, Tian Gan, and Liqiang Nie. 2020. LARA: Attribute-to-feature adversarial learning for new-item recommendation. In Proceedings of the 13th international conference on web search and data mining. 582–590.
- Sun et al. (2015) Jianshan Sun, Gang Wang, Xusen Cheng, and Yelin Fu. 2015. Mining affective text to improve social media item recommendation. Information Processing & Management 51, 4 (2015), 444–457.
- Sun et al. (2021) Mingze Sun, Daiyue Xue, Weipeng Wang, Qifu Hu, and Jianping Yu. 2021. Group-based deep transfer learning with mixed gate control for cross-domain recommendation. In 2021 International Joint Conference on Neural Networks (IJCNN). IEEE, 1–8.
- Tang et al. (2024b) Hao Tang, Guoshuai Zhao, Jing Gao, and Xueming Qian. 2024b. Personalized Representation With Contrastive Loss for Recommendation Systems. IEEE Trans. Multim. 26 (2024), 2419–2429. https://doi.org/10.1109/TMM.2023.3295740
- Tang et al. (2024a) Jiakai Tang, Sunhao Dai, Zexu Sun, Xu Chen, Jun Xu, Wenhui Yu, Lantao Hu, Peng Jiang, and Han Li. 2024a. Towards Robust Recommendation via Decision Boundary-aware Graph Contrastive Learning. In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. 2854–2865.
- Tao et al. (2022) Zhulin Tao, Xiaohao Liu, Yewei Xia, Xiang Wang, Lifang Yang, Xianglin Huang, and Tat-Seng Chua. 2022. Self-supervised learning for multimedia recommendation. IEEE Transactions on Multimedia (2022).
- Tian et al. (2022) Changxin Tian, Zihan Lin, Shuqing Bian, Jinpeng Wang, and Wayne Xin Zhao. 2022. Temporal Contrastive Pre-Training for Sequential Recommendation. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management. 1925–1934.
- Togashi et al. (2021) Riku Togashi, Mayu Otani, and Shin’ichi Satoh. 2021. Alleviating cold-start problems in recommendation through pseudo-labelling over knowledge graph. In Proceedings of the 14th ACM international conference on web search and data mining. 931–939.
- Tran et al. (2019) Thanh Tran, Renee Sweeney, and Kyumin Lee. 2019. Adversarial mahalanobis distance-based attentive song recommender for automatic playlist continuation. In Proceedings of the 42nd International ACM SIGIR Conference on Research and Development in Information Retrieval. 245–254.
- Wang et al. (2021b) Can Wang, Jiawei Chen, Sheng Zhou, Qihao Shi, Yan Feng, and Chun Chen. 2021b. SamWalker++: recommendation with informative sampling strategy. IEEE Transactions on Knowledge and Data Engineering (2021).
- Wang et al. (2023a) Chenyang Wang, Weizhi Ma, Chong Chen, Min Zhang, Yiqun Liu, and Shaoping Ma. 2023a. Sequential recommendation with multiple contrast signals. ACM Transactions on Information Systems 41, 1 (2023), 1–27.
- Wang et al. (2018a) Hongwei Wang, Jia Wang, Jialin Wang, Miao Zhao, Weinan Zhang, Fuzheng Zhang, Xing Xie, and Minyi Guo. 2018a. Graphgan: Graph representation learning with generative adversarial nets. In Proceedings of the AAAI conference on artificial intelligence, Vol. 32.
- Wang et al. (2019) Hongwei Wang, Fuzheng Zhang, Miao Zhao, Wenjie Li, Xing Xie, and Minyi Guo. 2019. Multi-task feature learning for knowledge graph enhanced recommendation. In The world wide web conference. 2000–2010.
- Wang et al. (2023b) Jiayin Wang, Weizhi Ma, Chumeng Jiang, Min Zhang, Yuan Zhang, Biao Li, and Peng Jiang. 2023b. Measuring Item Global Residual Value for Fair Recommendation. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval. 269–278.
- Wang et al. (2017b) Jun Wang, Lantao Yu, Weinan Zhang, Yu Gong, Yinghui Xu, Benyou Wang, Peng Zhang, and Dell Zhang. 2017b. Irgan: A minimax game for unifying generative and discriminative information retrieval models. In Proceedings of the 40th International ACM SIGIR conference on Research and Development in Information Retrieval. 515–524.
- Wang et al. (2017a) Quan Wang, Zhendong Mao, Bin Wang, and Li Guo. 2017a. Knowledge graph embedding: A survey of approaches and applications. IEEE Transactions on Knowledge and Data Engineering 29, 12 (2017), 2724–2743.
- Wang et al. (2018b) Qinyong Wang, Hongzhi Yin, Zhiting Hu, Defu Lian, Hao Wang, and Zi Huang. 2018b. Neural memory streaming recommender networks with adversarial training. In Proceedings of the 24th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining. 2467–2475.
- Wang et al. (2021c) Wenjie Wang, Fuli Feng, Xiangnan He, Liqiang Nie, and Tat-Seng Chua. 2021c. Denoising implicit feedback for recommendation. In Proceedings of the 14th ACM international conference on web search and data mining. 373–381.
- Wang et al. (2021a) Xin Wang, Yudong Chen, and Wenwu Zhu. 2021a. A survey on curriculum learning. IEEE Transactions on Pattern Analysis and Machine Intelligence 44, 9 (2021), 4555–4576.
- Wang et al. (2021d) Xi Wang, Iadh Ounis, and Craig Macdonald. 2021d. Leveraging review properties for effective recommendation. In Proceedings of the Web Conference 2021. 2209–2219.
- Wang et al. (2020) Xiang Wang, Yaokun Xu, Xiangnan He, Yixin Cao, Meng Wang, and Tat-Seng Chua. 2020. Reinforced negative sampling over knowledge graph for recommendation. In Proceedings of the web conference 2020. 99–109.
- Wang et al. (2024a) Yidan Wang, Zhaochun Ren, Weiwei Sun, Jiyuan Yang, Zhixiang Liang, Xin Chen, Ruobing Xie, Su Yan, Xu Zhang, Pengjie Ren, et al. 2024a. Enhanced generative recommendation via content and collaboration integration. arXiv preprint arXiv:2403.18480 (2024).
- Wang et al. (2022) Yu Wang, Hengrui Zhang, Zhiwei Liu, Liangwei Yang, and Philip S Yu. 2022. ContrastVAE: Contrastive Variational AutoEncoder for Sequential Recommendation. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management. 2056–2066.
- Wang et al. (2024b) Zihan Wang, Yonghui Yang, Le Wu, Richang Hong, and Meng Wang. 2024b. Making Non-overlapping Matters: An Unsupervised Alignment enhanced Cross-Domain Cold-Start Recommendation. IEEE Transactions on Knowledge and Data Engineering (2024).
- Wei et al. (2022) Moshi Wei, Nima Shiri Harzevili, Yuchao Huang, Junjie Wang, and Song Wang. 2022. CLEAR: Contrastive Learning for API Recommendation. In 2022 IEEE/ACM 44th International Conference on Software Engineering (ICSE).
- Wei et al. (2021) Yinwei Wei, Xiang Wang, Qi Li, Liqiang Nie, Yan Li, Xuanping Li, and Tat-Seng Chua. 2021. Contrastive learning for cold-start recommendation. In Proceedings of the 29th ACM International Conference on Multimedia. 5382–5390.
- Wu et al. (2023a) Bin Wu, Xiangnan He, Le Wu, Xue Zhang, and Yangdong Ye. 2023a. Graph-Augmented Co-Attention Model for Socio-Sequential Recommendation. IEEE Trans. Syst. Man Cybern. Syst. 53, 7 (2023), 4039–4051. https://doi.org/10.1109/TSMC.2023.3242308
- Wu et al. (2023b) Bin Wu, Xiangnan He, Qi Zhang, Meng Wang, and Yangdong Ye. 2023b. GCRec: Graph-Augmented Capsule Network for Next-Item Recommendation. IEEE Trans. Neural Networks Learn. Syst. 34, 12 (2023), 10164–10177. https://doi.org/10.1109/TNNLS.2022.3164982
- Wu and Ye (2020) Bin Wu and Yangdong Ye. 2020. BSPR: Basket-sensitive personalized ranking for product recommendation. Inf. Sci. 541 (2020), 185–206. https://doi.org/10.1016/J.INS.2020.06.046
- Wu et al. (2023c) Bin Wu, Lihong Zhong, and Yangdong Ye. 2023c. Graph-Augmented Social Translation Model for Next-Item Recommendation. IEEE Trans. Ind. Informatics 19, 11 (2023), 10913–10922. https://doi.org/10.1109/TII.2023.3242809
- Wu et al. (2019c) Chuhan Wu, Fangzhao Wu, Mingxiao An, Yongfeng Huang, and Xing Xie. 2019c. Neural news recommendation with topic-aware news representation. In Proceedings of the 57th Annual meeting of the association for computational linguistics. 1154–1159.
- Wu et al. (2021b) Chuhan Wu, Fangzhao Wu, Tao Qi, and Yongfeng Huang. 2021b. Empowering news recommendation with pre-trained language models. In Proceedings of the 44th International ACM SIGIR Conference on Research and Development in Information Retrieval.
- Wu et al. (2022b) Chuhan Wu, Fangzhao Wu, Tao Qi, Qi Liu, Xuan Tian, Jie Li, Wei He, Yongfeng Huang, and Xing Xie. 2022b. Feedrec: News feed recommendation with various user feedbacks. In Proceedings of the ACM Web Conference 2022. 2088–2097.
- Wu et al. (2022a) Haolun Wu, Chen Ma, Yingxue Zhang, Xue Liu, Ruiming Tang, and Mark Coates. 2022a. Adapting Triplet Importance of Implicit Feedback for Personalized Recommendation. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management. 2148–2157.
- Wu et al. (2024a) Junkang Wu, Jiawei Chen, Jiancan Wu, Wentao Shi, Jizhi Zhang, and Xiang Wang. 2024a. BSL: Understanding and Improving Softmax Loss for Recommendation. In 40th IEEE International Conference on Data Engineering, ICDE 2024, Utrecht, The Netherlands, May 13-16, 2024. IEEE, 816–830. https://doi.org/10.1109/ICDE60146.2024.00068
- Wu et al. (2021a) Jiancan Wu, Xiang Wang, Fuli Feng, Xiangnan He, Liang Chen, Jianxun Lian, and Xing Xie. 2021a. Self-supervised graph learning for recommendation. In Proceedings of the 44th international ACM SIGIR conference on research and development in information retrieval. 726–735.
- Wu et al. (2024b) Jiancan Wu, Xiang Wang, Xingyu Gao, Jiawei Chen, Hongcheng Fu, and Tianyu Qiu. 2024b. On the Effectiveness of Sampled Softmax Loss for Item Recommendation. ACM Trans. Inf. Syst. 42, 4 (2024), 98:1–98:26. https://doi.org/10.1145/3637061
- Wu et al. (2019a) Le Wu, Lei Chen, Yonghui Yang, Richang Hong, Yong Ge, Xing Xie, and Meng Wang. 2019a. Personalized multimedia item and key frame recommendation. In Proceedings of the 28th International Joint Conference on Artificial Intelligence. 1431–1437.
- Wu et al. (2024c) Likang Wu, Zhi Zheng, Zhaopeng Qiu, Hao Wang, Hongchao Gu, Tingjia Shen, Chuan Qin, Chen Zhu, Hengshu Zhu, Qi Liu, et al. 2024c. A survey on large language models for recommendation. World Wide Web 27, 5 (2024), 60.
- Wu et al. (2019b) Qitian Wu, Yirui Gao, Xiaofeng Gao, Paul Weng, and Guihai Chen. 2019b. Dual sequential prediction models linking sequential recommendation and information dissemination. In Proceedings of the 25th ACM SIGKDD international conference on knowledge discovery & data mining. 447–457.
- Wu et al. (2024d) Xuansheng Wu, Huachi Zhou, Yucheng Shi, Wenlin Yao, Xiao Huang, and Ninghao Liu. 2024d. Could Small Language Models Serve as Recommenders? Towards Data-centric Cold-start Recommendation. In Proceedings of the ACM on Web Conference. 3566–3575.
- Wu et al. (2019d) Yongkai Wu, Lu Zhang, and Xintao Wu. 2019d. Counterfactual fairness: Unidentification, bound and algorithm. In Proceedings of the twenty-eighth international joint conference on Artificial Intelligence.
- Xi et al. (2024) Yunjia Xi, Weiwen Liu, Jianghao Lin, Xiaoling Cai, Hong Zhu, Jieming Zhu, Bo Chen, Ruiming Tang, Weinan Zhang, and Yong Yu. 2024. Towards open-world recommendation with knowledge augmentation from large language models. In Proceedings of the 18th ACM Conference on Recommender Systems. 12–22.
- Xia et al. (2022) Lianghao Xia, Chao Huang, Yong Xu, Peng Dai, and Liefeng Bo. 2022. Multi-Behavior Graph Neural Networks for Recommender System. IEEE Transactions on Neural Networks and Learning Systems (2022).
- Xia et al. (2021) Lianghao Xia, Chao Huang, Yong Xu, Peng Dai, Xiyue Zhang, Hongsheng Yang, Jian Pei, and Liefeng Bo. 2021. Knowledge-enhanced hierarchical graph transformer network for multi-behavior recommendation. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 35. 4486–4493.
- Xie et al. (2021) Ruobing Xie, Cheng Ling, Yalong Wang, Rui Wang, Feng Xia, and Leyu Lin. 2021. Deep feedback network for recommendation. In Proceedings of the Twenty-Ninth International Conference on International Joint Conferences on Artificial Intelligence. 2519–2525.
- Xie et al. (2022b) Ruobing Xie, Shaoliang Zhang, Rui Wang, Feng Xia, and Leyu Lin. 2022b. A peep into the future: Adversarial future encoding in recommendation. In Proceedings of the Fifteenth ACM International Conference on Web Search and Data Mining. 1177–1185.
- Xie et al. (2022a) Xu Xie, Fei Sun, Zhaoyang Liu, Shiwen Wu, Jinyang Gao, Jiandong Zhang, Bolin Ding, and Bin Cui. 2022a. Contrastive learning for sequential recommendation. In 2022 IEEE 38th international conference on data engineering (ICDE). IEEE, 1259–1273.
- Xu et al. (2021) Kerui Xu, Jingxuan Yang, Jun Xu, Sheng Gao, Jun Guo, and Ji-Rong Wen. 2021. Adapting user preference to online feedback in multi-round conversational recommendation. In Proceedings of the 14th ACM international conference on web search and data mining. 364–372.
- Xu et al. (2022) Lanling Xu, Jianxun Lian, Wayne Xin Zhao, Ming Gong, Linjun Shou, Daxin Jiang, Xing Xie, and Ji-Rong Wen. 2022. Negative sampling for contrastive representation learning: A review. arXiv preprint arXiv:2206.00212 (2022).
- Yang et al. (2020b) Carl Yang, Yuxin Xiao, Yu Zhang, Yizhou Sun, and Jiawei Han. 2020b. Heterogeneous network representation learning: A unified framework with survey and benchmark. IEEE Transactions on Knowledge and Data Engineering 34, 10 (2020), 4854–4873.
- Yang et al. (2023a) Fan Yang, Zheng Chen, Ziyan Jiang, Eunah Cho, Xiaojiang Huang, and Yanbin Lu. 2023a. PALR: Personalization Aware LLMs for Recommendation. (2023). http://arxiv.org/abs/2305.07622
- Yang et al. (2021a) Haoran Yang, Hongxu Chen, Lin Li, S Yu Philip, and Guandong Xu. 2021a. Hyper meta-path contrastive learning for multi-behavior recommendation. In 2021 IEEE International Conference on Data Mining (ICDM). IEEE, 787–796.
- Yang et al. (2020c) Ji Yang, Xinyang Yi, Derek Zhiyuan Cheng, Lichan Hong, Yang Li, Simon Xiaoming Wang, Taibai Xu, and Ed H Chi. 2020c. Mixed negative sampling for learning two-tower neural networks in recommendations. In Companion Proceedings of the Web Conference 2020. 441–447.
- Yang et al. (2023b) Liangwei Yang, Shengjie Wang, Yunzhe Tao, Jiankai Sun, Xiaolong Liu, Philip S Yu, and Taiqing Wang. 2023b. Dgrec: Graph neural network for recommendation with diversified embedding generation. In Proceedings of the sixteenth ACM international conference on web search and data mining. 661–669.
- Yang et al. (2021b) Mengyue Yang, Quanyu Dai, Zhenhua Dong, Xu Chen, Xiuqiang He, and Jun Wang. 2021b. Top-n recommendation with counterfactual user preference simulation. In Proceedings of the 30th ACM International Conference on Information & Knowledge Management. 2342–2351.
- Yang et al. (2021c) Sean Bin Yang, Chenjuan Guo, Jilin Hu, Jian Tang, and Bin Yang. 2021c. Unsupervised Path Representation Learning with Curriculum Negative Sampling. Proceedings of International Joint Conference on Artificial Intelligence (IJCAI) (2021).
- Yang et al. (2022a) Yuhao Yang, Chao Huang, Lianghao Xia, and Chenliang Li. 2022a. Knowledge graph contrastive learning for recommendation. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval. 1434–1443.
- Yang et al. (2022b) Yuhao Yang, Chao Huang, Lianghao Xia, Yuxuan Liang, Yanwei Yu, and Chenliang Li. 2022b. Multi-behavior hypergraph-enhanced transformer for sequential recommendation. In Proceedings of the 28th ACM SIGKDD conference on knowledge discovery and data mining. 2263–2274.
- Yang et al. (2021d) Yonghui Yang, Le Wu, Richang Hong, Kun Zhang, and Meng Wang. 2021d. Enhanced graph learning for collaborative filtering via mutual information maximization. In Proceedings of the 44th International ACM SIGIR Conference on Research and Development in Information Retrieval. 71–80.
- Yang et al. (2024) Yonghui Yang, Le Wu, Zihan Wang, Zhuangzhuang He, Richang Hong, and Meng Wang. 2024. Graph bottlenecked social recommendation. In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. 3853–3862.
- Yang et al. (2023d) Yonghui Yang, Le Wu, Kun Zhang, Richang Hong, Hailin Zhou, Zhiqiang Zhang, Jun Zhou, and Meng Wang. 2023d. Hyperbolic graph learning for social recommendation. IEEE Transactions on Knowledge and Data Engineering 36, 12 (2023), 8488–8501.
- Yang et al. (2023c) Yonghui Yang, Zhengwei Wu, Le Wu, Kun Zhang, Richang Hong, Zhiqiang Zhang, Jun Zhou, and Meng Wang. 2023c. Generative-contrastive graph learning for recommendation. In Proceedings of the 46th international ACM SIGIR Conference on Research and Development in Information Retrieval. 1117–1126.
- Yang et al. (2020a) Zhen Yang, Ming Ding, Chang Zhou, Hongxia Yang, Jingren Zhou, and Jie Tang. 2020a. Understanding negative sampling in graph representation learning. In Proceedings of the 26th ACM SIGKDD international conference on knowledge discovery & data mining. 1666–1676.
- Yin et al. (2019) Hongzhi Yin, Qinyong Wang, Kai Zheng, Zhixu Li, Jiali Yang, and Xiaofang Zhou. 2019. Social influence-based group representation learning for group recommendation. In 2019 IEEE 35th International Conference on Data Engineering (ICDE). IEEE, 566–577.
- Ying et al. (2018) Rex Ying, Ruining He, Kaifeng Chen, Pong Eksombatchai, William L Hamilton, and Jure Leskovec. 2018. Graph convolutional neural networks for web-scale recommender systems. In Proceedings of the 24th ACM SIGKDD international conference on knowledge discovery & data mining. 974–983.
- Yu et al. (2018) Junliang Yu, Min Gao, Jundong Li, Hongzhi Yin, and Huan Liu. 2018. Adaptive implicit friends identification over heterogeneous network for social recommendation. In Proceedings of the 27th ACM international conference on information and knowledge management. 357–366.
- Yu et al. (2020) Jiangxing Yu, Hong Zhu, Chih-Yao Chang, Xinhua Feng, Bowen Yuan, Xiuqiang He, and Zhenhua Dong. 2020. Influence function for unbiased recommendation. In Proceedings of the 43rd International ACM SIGIR Conference on Research and Development in Information Retrieval. 1929–1932.
- Zang et al. (2022) Tianzi Zang, Yanmin Zhu, Haobing Liu, Ruohan Zhang, and Jiadi Yu. 2022. A survey on cross-domain recommendation: taxonomies, methods, and future directions. ACM Transactions on Information Systems 41, 2 (2022), 1–39.
- Zhang et al. (2023b) An Zhang, Leheng Sheng, Zhibo Cai, Xiang Wang, and Tat-Seng Chua. 2023b. Empowering Collaborative Filtering with Principled Adversarial Contrastive Loss. In Advances in Neural Information Processing Systems 36: Annual Conference on Neural Information Processing Systems 2023, NeurIPS 2023, New Orleans, LA, USA, December 10 - 16, 2023. http://papers.nips.cc/paper_files/paper/2023/hash/13f1750b825659394a6499399e7637fc-Abstract-Conference.html
- Zhang et al. (2023a) Jizhi Zhang, Keqin Bao, Yang Zhang, Wenjie Wang, Fuli Feng, and Xiangnan He. 2023a. Is ChatGPT Fair for Recommendation? Evaluating Fairness in Large Language Model Recommendation. (2023). https://doi.org/10.1145/3604915.3608860
- Zhang et al. (2021a) Junwei Zhang, Min Gao, Junliang Yu, Lei Guo, Jundong Li, and Hongzhi Yin. 2021a. Double-scale self-supervised hypergraph learning for group recommendation. In Proceedings of the 30th ACM International Conference on Information & Knowledge Management. 2557–2567.
- Zhang et al. (2024a) Kepu Zhang, Teng Shi, Sunhao Dai, Xiao Zhang, Yinfeng Li, Jing Lu, Xiaoxue Zang, Yang Song, and Jun Xu. 2024a. SAQRec: Aligning Recommender Systems to User Satisfaction via Questionnaire Feedback. In Proceedings of the 33rd ACM International Conference on Information and Knowledge Management. 3165–3175.
- Zhang et al. (2021b) Shengyu Zhang, Dong Yao, Zhou Zhao, Tat-Seng Chua, and Fei Wu. 2021b. Causerec: Counterfactual user sequence synthesis for sequential recommendation. In Proceedings of the 44th International ACM SIGIR Conference on Research and Development in Information Retrieval. 367–377.
- Zhang et al. (2024b) Tao Zhang, Luwei Yang, Zhibo Xiao, Wen Jiang, and Wei Ning. 2024b. On Practical Diversified Recommendation with Controllable Category Diversity Framework. In Companion Proceedings of the ACM on Web Conference 2024. 255–263.
- Zhang et al. (2013) Weinan Zhang, Tianqi Chen, Jun Wang, and Yong Yu. 2013. Optimizing top-n collaborative filtering via dynamic negative item sampling. In Proceedings of the 36th international ACM SIGIR conference on Research and development in information retrieval.
- Zhao et al. (2019) Jun Zhao, Zhou Zhou, Ziyu Guan, Wei Zhao, Wei Ning, Guang Qiu, and Xiaofei He. 2019. Intentgc: a scalable graph convolution framework fusing heterogeneous information for recommendation. In Proceedings of the 25th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining. 2347–2357.
- Zhao et al. (2016) Wayne Xin Zhao, Jinpeng Wang, Yulan He, Ji-Rong Wen, Edward Y Chang, and Xiaoming Li. 2016. Mining product adopter information from online reviews for improving product recommendation. ACM Transactions on Knowledge Discovery from Data (TKDD) 10, 3 (2016), 1–23.
- Zhao et al. (2023) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al. 2023. A survey of large language models. arXiv preprint arXiv:2303.18223 (2023).
- Zhao et al. (2018) Xiangyu Zhao, Liang Zhang, Zhuoye Ding, Long Xia, Jiliang Tang, and Dawei Yin. 2018. Recommendations with negative feedback via pairwise deep reinforcement learning. In Proceedings of the 24th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining. 1040–1048.
- Zheng et al. (2018) Guanjie Zheng, Fuzheng Zhang, Zihan Zheng, Yang Xiang, Nicholas Jing Yuan, Xing Xie, and Zhenhui Li. 2018. DRN: A deep reinforcement learning framework for news recommendation. In Proceedings of the 2018 world wide web conference. 167–176.
- Zheng et al. (2022) Xiaolin Zheng, Jiajie Su, Weiming Liu, and Chaochao Chen. 2022. DDGHM: Dual Dynamic Graph with Hybrid Metric Training for Cross-Domain Sequential Recommendation. In Proceedings of the 30th ACM International Conference on Multimedia. 471–481.
- Zheng et al. (2021) Yu Zheng, Chen Gao, Xiang Li, Xiangnan He, Yong Li, and Depeng Jin. 2021. Disentangling user interest and conformity for recommendation with causal embedding. In Proceedings of the Web Conference 2021. 2980–2991.
- Zhou et al. (2018) Chang Zhou, Jinze Bai, Junshuai Song, Xiaofei Liu, Zhengchao Zhao, Xiusi Chen, and Jun Gao. 2018. Atrank: An attention-based user behavior modeling framework for recommendation. In Proceedings of the AAAI conference on artificial intelligence, Vol. 32.
- Zhou et al. (2020) Kun Zhou, Hui Wang, Wayne Xin Zhao, Yutao Zhu, Sirui Wang, Fuzheng Zhang, Zhongyuan Wang, and Ji-Rong Wen. 2020. S3-rec: Self-supervised learning for sequential recommendation with mutual information maximization. In Proceedings of the 29th ACM international conference on information & knowledge management. 1893–1902.
- Zhou and Shen (2023) Xin Zhou and Zhiqi Shen. 2023. A tale of two graphs: Freezing and denoising graph structures for multimodal recommendation. In Proceedings of the 31st ACM International Conference on Multimedia. 935–943.
- Zhou et al. (2021) Yao Zhou, Jianpeng Xu, Jun Wu, Zeinab Taghavi, Evren Korpeoglu, Kannan Achan, and Jingrui He. 2021. PURE: Positive-unlabeled recommendation with generative adversarial network. In Proceedings of the 27th ACM SIGKDD Conference on Knowledge Discovery & Data Mining. 2409–2419.
- Zhu et al. (2018) Han Zhu, Xiang Li, Pengye Zhang, Guozheng Li, Jie He, Han Li, and Kun Gai. 2018. Learning tree-based deep model for recommender systems. In Proceedings of the 24th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining. 1079–1088.
- Zhu et al. (2022) Qiannan Zhu, Haobo Zhang, Qing He, and Zhicheng Dou. 2022. A Gain-Tuning Dynamic Negative Sampler for Recommendation. In Proceedings of the ACM Web Conference 2022. 277–285.
- Zhuo et al. (2022) Jianhuan Zhuo, Qiannan Zhu, Yinliang Yue, and Yuhong Zhao. 2022. Learning explicit user interest boundary for recommendation. In Proceedings of the ACM Web Conference 2022. 193–202.
- Zou et al. (2024) Ding Zou, Wei Wei, Feida Zhu, Chuanyu Xu, Tao Zhang, and Chengfu Huo. 2024. Knowledge Enhanced Multi-intent Transformer Network for Recommendation. In Companion Proceedings of the ACM on Web Conference 2024. 1–9.